### v3

Abstract
In this paper, we tackle the dynamic mean-variance portfolio selection problem in a model-free manner, based on (generative) diffusion models. We propose using data sampled from the real model $\mathbb{P}$ (which is unknown) with limited size to train a generative model $\mathbb{Q}$ (from which we can easily and adequately sample). With adaptive training and sampling methods that are tailor-made for time-series data, we obtain quantification bounds between $\mathbb{P}$ and $\mathbb{Q}$ in terms of the adapted Wasserstein metric $\mathcal{A}\mathcal{W}_{2}$. Importantly, the proposed adapted sampling method also facilitates conditional sampling. In the second part of this paper, we provide the stability of the mean-variance portfolio optimization problems in $\mathcal{A}\mathcal{W}_{2}$. Then, combined with the error bounds and the stability result, we propose a policy gradient algorithm based on the generative environment, in which our innovative adapted sampling method provides approximate scenario generators. We illustrate the performance of our algorithm on both simulated and real data. For real data, the algorithm based on the generative environment produces portfolios that beat several important baselines, including the Markowitz portfolio, the equal weight (naive) portfolio, and S&P 500.
Subject Classification: dynamic programming: optimal control, finance: portfolio, computer science: artificial intelligence.
1 Introduction
Portfolio selection, a crucial and fundamental task in financial engineering, has been a subject of both theoretical and practical interest since the classical work Markowitz (1952). With the ability to account for intertemporal hedging and rebalancing in a non-stationary market, formulating portfolio selection as dynamic multi-period problems has become increasingly appealing; see Mossin (1968), Merton (1971) for early works. Researchers have explored this problem under various specifications, such as continuous-time dynamic mean-variance problems, where return models are prescribed by some stochastic differential equations Zhou and Li (2000), or discrete-time problems with independent intertemporal returns Li and Ng (2000). However, due to incompleteness of the information, estimation errors, or simply the infeasibility of fully modeling market uncertainty, we lack access to the real model, as opposed to model-based approaches. Instead, we are restricted to a limited set of financial data samples. Consequently, investigating portfolio selection problems in a model-free manner has been a long-standing topic in financial engineering; see discussions of ‘universal portfolio’ Cover (1991), stochastic portfolio theory Karatzas and Ruf (2017), Cuchiero et al. (2019), and the application of machine learning Ban et al. (2018), Gu et al. (2020). However, dynamic portfolio selection problems are rather involved, even with model-based approaches, due to their inherent intertemporal structure Campbell and Viceira (1999), high dimensionality Brandt et al. (2005), and time inconsistency Basak and Chabakauri (2010), Björk et al. (2014). Therefore, model-free dynamic portfolio selection problems are worth exploring.
To address this challenge, we propose to leverage score-based diffusion models. These models can generate additional samples that exhibit distributional properties similar to the original data. Specifically, when the model $s^{1:T}\sim\mathbb{P}$ is unknown, we propose a data-driven approach to train a model $\mathbb{Q}$ that closely resembles $\mathbb{P}$. We utilize the model $\mathbb{Q}$ as an oracle in the subsequent policy gradient algorithm: we sample a large number of paths $\tilde{s}^{1:T}\sim\mathbb{Q}$ and train a reinforcement learning (RL) agent using these simulated paths. Theoretically, we show that our method achieves near-optimality in terms of Markowitz’s mean-variance criterion, thanks to two key contributions of the present paper: (1) a quantification bound between $\mathbb{P}$ and $\mathbb{Q}$, and (2) the model stability. From an engineering perspective, the approximated model $\mathbb{Q}$ consists of a score network $s_{\theta}$, which is the core of the score-matching technique used in generative models, and a Recurrent Neural Network (RNN) encoder $R_{\theta}$, which encodes the current state of the market. These two networks are trained simultaneously using historical price data. The reinforcement learning agent is trained on the generated data from $\mathbb{Q}$, which are produced by evaluating $s_{\theta}$ and $R_{\theta}$ in an adaptive manner. Finally, the agent, fed with the encoded (and updated) market state, outputs portfolio selection actions whose performance is evaluated through real-world experiments. For a summary of the theoretical results, see Figure 1, and for the workflow when implementing the proposed solution for model-free dynamic portfolio selection, refer to Figure 2.
[…]
Figure 1: Model-free dynamic MV portfolio selection: theory
[…]
Figure 2: Model-free dynamic MV portfolio selection: implementation
1.1 Background, main results and contributions
The score matching technique, which has been a powerful tool in statistics since the fundamental work Hyvärinen (2005), has also been incorporated into generative models in various works, including Sohl-Dickstein et al. (2015), Song and Ermon (2019), and others. Specifically, Song et al. (2021b) unifies the concept of score-based generative models and diffusion probabilistic models (Ho et al. (2020)) into a single framework, known as score-based diffusion models, which we adopt in this work.
Traditionally, a score‑based diffusion model is defined through a forward stochastic differential equation (SDE) whose initial condition is the data distribution. A popular choice—adopted in this paper—is the variance‑preserving SDE (VP‑SDE). Specifically, let $p_{\rm data}$ be a probability distribution on $\mathbb{R}^{d}$ from which the data are drawn. The VP‑SDE is given by
$\begin{cases}&\mathrm{d}X_{\tau}=-\tfrac{1}{2}\,\beta(\tau)\,X_{\tau}\,\mathrm{d}\tau+\sqrt{\beta(\tau)}\,\mathrm{d}B_{\tau},\\ &X_{0}\sim p_{\rm data}\end{cases}$ (1.1)
Here $\beta(\tau)=\beta_{\min}+(\beta_{\max}-\beta_{\min}),\tau$ is a prescribed drift schedule. Although $p_{\rm data}$ is unknown, the SDE in (1.1) converges to the invariant distribution $\mathcal{N}(0,I_{d\times d})$ as $\tau\to\infty$. Diffusion models therefore consider the following reverse‑time SDE:
$\begin{cases}&dY_{\tau}=\bigg{(}\frac{1}{2}\beta(\mathcal{T}-\tau)Y_{\tau}+\beta(\mathcal{T}-\tau)\nabla\log p(\mathcal{T}-\tau,Y_{\tau})\bigg{)}\mathrm{d}\tau+\sqrt{\beta(\mathcal{T}-\tau)}\mathrm{d}\bar{B}_{\tau},\\ &Y_{0}\sim\mathcal{N}(0,I_{d\times d}),\end{cases}$ (1.2)
where $p(\tau,\cdot)$ denotes the density of $X_{\tau}$. Because $p_{\rm data}$ is unknown, so is $p(\tau,\cdot)$. Under the score‑matching theory, an accurate estimate of the score $\nabla\log p$ guarantees that, for sufficiently large $\mathcal{T}$, the law of $Y_{\mathcal{T}}$ approximates $p_{\rm data}$ under various metrics; see Tang and Zhao (2025) for a comprehensive survey and Gao et al. (2025) for recent Wasserstein bounds. Score-based diffusion models have found numerous applications across a wide range of areas, including image and video generation (Rombach et al. (2022), Melnik et al. (2024)), molecule simulation Hoogeboom et al. (2022) and health Abramson et al. (2024). A very recent work Chen et al. (2025) has also specifically applied the score-based diffusion model to financial data generation, exploiting low-dimensional structures in high-dimensional return data.
Nevertheless, because we are interested in dynamic decision-making problems, the data take the form of time-series: $X=(X^{1},X^{2},\cdots,X^{T})=X^{1:T}$. For instance, $T=12$ and each $X^{t}$ is the monthly price data, then $X$ is a one-year price path. Although score-based diffusion models have been applied to time-series prediction (see Lin et al. (2024) for a survey), to the best of our knowledge, there is still no theoretical result that establishes foundation for time-series data generation. Treating the state space naively as the enlarged Euclidean space $\mathbb{R}^{dT}$ is unsatisfactory: it overlooks the intertemporal structure and the non-anticipation requirements inherent in dynamic problems. Indeed, existing results show that, for dynamic settings, value functions are unstable under standard probabilistic metrics such as Wasserstein; see, e.g., Backhoff-Veraguas et al. (2020). This observation motivates the use of the adapted Wasserstein metric, $\mathcal{A}\mathcal{W}_{2}$, which is better suited to time-series data and dynamic tasks; see Backhoff-Veraguas et al. (2020) for definitions. Furthermore, a dynamic control problem requires sampling not only from the joint distribution of $\mathbb{P}$ but also from the conditional distribution $\mathbb{P}(\cdot|X^{1:t}=x^{1:t})$ given observations up to time $t$. Such conditional sampling is indispensable for our subsequent policy-gradient algorithm (i.e., for training the RL agent). Traditional score-based diffusion models, however, do not support this functionality. Consequently, we must develop new training and sampling procedures that (i) admit theoretical score-matching error bounds under the adapted Wasserstein metric and (ii) enable conditional sampling for any observation input.
Our first set of results meets these two requirements. For notational convenience, we fix $\beta\equiv 2$ in (1.1) and, for each $t=1,2,\dots,T-1$, consider the following conditional VP-SDE:
$\begin{cases}&\mathrm{d}X^{t+1}_{\tau}=-X^{t+1}_{\tau}\mathrm{d}\tau+\sqrt{2}\mathrm{d}B^{t+1}_{\tau},\\ &X^{t+1}_{0}\sim\mathbb{P}_{x^{1:t}},\end{cases}$ (1.3)
where $\mathbb{P}_{x^{1:t}}$ denotes the law of $X^{t+1}$ conditional on $X^{1:t}=x^{1:t}$. This formulation allows us to define the conditional score function $s^{t+1}(\tau,x^{1:t},x)=\nabla_{x}\log p_{t+1}(\tau,x^{1:t},x)$, where $p_{t+1}(\tau,x^{1:t},\cdot)$ is the density of $X^{t+1}_{\tau}$ in (1.3). Assume that we have an approximator $\hat{s}^{t+1}$ that is close to $s$ in a suitable $L^{2}$ sense. We generate a new trajectory $y^{1:T}$ adaptively: to obtain $y^{1}$ we start from pure noise and run the reverse VP-SDE (1.2), setting $y^{1}=Y_{\mathcal{T}}$; to obtain $y^{t+1}$ we use the previously generated path $y^{1:t}$ as conditional input and run (1.2) with the score replaced by $\hat{s}^{t+1}(\cdot,y^{1:t},\cdot)$. See Algorithm 1 for more details. Under appropriate technical conditions, we obtain the following bound:
Theorem (Theorem 2.3 in the main context).
Denote by $\mathbb{Q}^{\mathcal{T}}$ the distribution of $y^{1:T}$. Then $\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq E(\mathcal{T},\varepsilon_{\rm score})$, where $\varepsilon_{\rm score}$ is the score-matching error between $\hat{s}$ and $s$. Moreover, $E(\mathcal{T},\varepsilon_{\rm score})$ has an explicit form and can be made arbitrarily small by taking $\mathcal{T}$ sufficiently large and $\varepsilon_{\rm score}$ sufficiently small.
Crucially, the adaptive sampling scheme also enables conditional sampling: given an observation $x^{1:t}$, we simply treat it as the conditional input and use $\hat{s}^{t+1}(\cdot,x^{1:t},\cdot)$ in the reverse SDE. The proof of these error bounds rests on three key technical ingredients: (i) classical total variation bounds between conditional distributions, (ii) a new set of technical assumptions on the (joint) data distribution, and (iii) explicit tail estimates for the SDE. See Section 2 and the appendices for details.
While we can now generate new samples from $\mathbb{Q}$, several steps remain before we can fully address the dynamic portfolio selection problem. First, because we now work under the surrogate model $\mathbb{Q}$, we must establish model stability: even if $\mathbb{P}$ and $\mathbb{Q}$ are close and we can solve the mean–variance problem under $\mathbb{Q}$, does the optimizer obtained from $\mathbb{Q}$ remain near‑optimal for the original model $\mathbb{P}$? We answer this question in the affirmative:
Theorem (Theorem 3.8 and Corollary 3.9 in the main context).
Denote by $v(\mathbb{P})$ and $v(\mathbb{Q})$ the optimal values of the mean–variance problem under $\mathbb{P}$ and $\mathbb{Q}$, respectively. Then, under appropriate technical conditions, $|v(\mathbb{P})-v(\mathbb{Q})|\leq C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})$.
The final, yet still important and challenging, component of our framework is to efficiently solve the mean–variance problem under $\mathbb{Q}$. Although we can now sample from $\mathbb{Q}$, the distribution is implicit, so numerical methods are required. Drawing inspiration from recent advances in reinforcement learning (RL) within financial engineering (see Hambly et al. (2023) for a survey), we adopt a deep RL approach and propose a policy‑gradient algorithm (Algorithm 4). This is not a straightforward application of classical policy‑gradient methods because the problem possesses two distinctive features. First, portfolio selection is constrained by data inadequacy and lacks an episodic interaction structure: exploration cannot rely on repeatedly resetting the environment unless one uses a large archive of historical data, which may fail to reflect current market conditions. Instead, we generate fresh trajectories from $\mathbb{Q}$ using the most recent market information, thereby promoting relevant exploration. Second, standard policy‑gradient theory hinges on the dynamic‑programming principle (or the Bellman optimality equation), whereas the mean–variance objective is known to violate dynamic programming Basak and Chabakauri (2010). To overcome this obstacle, we employ a duality result from Labbé and Heunis (2007), which reduces the primal problem to a dynamic quadratic‑hedging formulation. Consequently, our policy‑gradient algorithm effectively solves the dual problem, and the resultant portfolio policy is obtained by plugging the optimal multiplier back into the duality relation. We further remark that this dual representation is instrumental in the proof of model stability; see Section 3 for details.
To demonstrate the practicality of the proposed approach, we implement the complete workflow shown in Figure 2. During the adaptive training and sampling of $\mathbb{Q}$, the state space $\mathbb{R}^{dt}$ expands sequentially for $t=1,2,\dots,T$, resulting in roughly $dT^{2}$ degrees of complexity. To curb this high dimensionality, we embed an RNN encoder that compresses $x^{1:t}$ into a feature space whose dimension is independent of $t$; see Rasul et al. (2021), Yan et al. (2021). In practice, every conditional input $x^{1:t}$ is replaced by its RNN encoding, and the encoder is trained jointly with the score network $s_{\theta}$. Section 4.3 provides a comprehensive empirical evaluation in both synthetic and real-world settings. In the synthetic experiment, where the true sampler is available, we present training logs showing that the policy gradient learns towards a model-based benchmark. In the real-world study, our strategy produces portfolios that beat the S&P 500, an equal-weight portfolio, and a historical-parameter Markowitz portfolio. We also investigate another diffusion-based strategy that constructs a Markowitz portfolio using parameters estimated from data generated conditional on the current market state. Although this diffusion-based Markowitz strategy is less stable than the policy-gradient approach, it still surpasses the benchmarks and, for certain risk-aversion levels, achieves the best overall performance. These findings highlight the effectiveness of the new training and sampling procedure for time-series data.
To conclude, the contributions of this work are threefold:
  1. 1.
    We extend the score‑based diffusion‑model framework to time‑series data. Using a novel adaptive training and sampling scheme, we establish theoretical error bounds under the adapted Wasserstein metric. The framework also supports conditional sampling, which is crucial for solving model‑free portfolio‑selection problems and constructing high‑performance portfolio strategies;
  2. 2.
    We prove a model‑stability result in the adapted Wasserstein metric. Because we focus on the mean–variance problem, the proof relies on a duality argument and a detailed analysis of the associated quadratic‑hedging problem;
  3. 3.
    We propose a policy‑gradient algorithm that operates in a surrogate environment generated by the diffusion model and exploits the dual formulation and dynamic programming. The algorithm, together with a diffusion‑based Markowitz strategy, outperforms several benchmarks. Experimental results validate both the adaptive training-sampling procedure and the effectiveness of conditional sampling.
1.2 Related works
The use of simulated price paths to inform investment and corporate‐finance decisions predates modern portfolio theory. Hertz (1964) first popularized Monte Carlo simulation for capital‐budgeting analysis, and Boyle (1977) introduced path‐wise Monte Carlo methods to option valuation. During the 1970s–1990s, more sophisticated parametric dynamics, jumping diffusion Merton (1976), stochastic volatility Heston (1993), and SV with jumps Bates (1996) - made large-scale pathwise risk analysis computationally tractable.
Monte Carlo simulation entered mainstream risk management when the RiskMetrics Technical Document popularized a scalable Value-at-Risk (VaR) framework J. P. Morgan (1996). Subsequent regulatory amendments (Basel II, 2004 ff.) accelerated the adoption of simulation as a predominant stress-testing device. Rockafellar and Uryasev (2000) then recast tail control as an optimisation of Conditional VaR (CVaR), turning scenario-weighted losses into a direct portfolio-construction objective.
Simulation also became a tool for taming estimation error. Michaud (1998)’s Resampled Efficient Frontier averages optimal weights across thousands of bootstrap resamples, trading a slight bias for a large variance reduction; the statistical foundations rest on Efron and Tibshirani (1994). Variance reduction techniques, such as antithetic variates, control variates, stratified sampling, are exhaustively cataloged by Glasserman (2004) and remain standard in parametric Monte Carlo.
A complementary line of work fits structured return models, simulates fresh paths, and optimises portfolios on those synthetic scenarios. For example, Paolella et al. (2021) feed non-Gaussian orthogonal-GARCH simulations into a mean–CVaR programme with transaction-cost penalties; Sahamkhadam et al. (2018) combine ARMA–GARCH marginals with EVT-copula tails to build minimum-CVaR frontiers; Ang and Bekaert (2002) Monte-Carlo a hidden-Markov VAR to capture regime shifts in international equity returns; Guastaroba et al. (2009) show multivariate-GARCH scenarios dominate historical or bootstrap windows inside a mean–CVaR model; and Lim (2004) derive closed-form mean–variance weights under stochastic-parameter diffusions, validating them via large-scale wealth simulations.
Across parametric simulation, and resampling, classical research progressively broadens robustness but still relies on explicit distributional or structural assumptions. Because these methods remain only partially data-driven, they can falter when markets exhibit complex, non-stationary joint tails or evolving dependence structures.
As asset universes became high-dimensional and tail dependencies more intricate, researchers began exploring fully data-driven generators. Early adversarial approaches such as TimeGAN Yoon et al. (2019) and QuantGAN Wiese et al. (2020) demonstrated that generative adversarial networks (GANs) can reproduce heavy tails, volatility clustering and cross-sectional dependence. Specialised variants soon followed: CorrGAN targets realistic correlation matrices for risk-parity and stress-testing tasks Marti (2020); Sig-Wasserstein GAN leverages the path-signature $W_{1}$ metric for conditional sampling that adapts to prevailing market states Liao et al. (2024); VolGAN specializes in generating implied volatility surfaces and is applied in constructing hedging portfolio Vuletić and Cont (2024), Cont and Vuletić (2025); and Tail-GAN, trained with a VaR/ES-based objective that explicitly focuses learning on severe left-tail outcomes, enriches downside stress scenarios beyond what standard GANs provide Cont et al. (2025). These models supply high-resolution scenario sets that far outstrip traditional resampling in capturing dependence structures, regime shifts and tail behaviour, but they remain prone to training instability Mescheder et al. (2018) and offer no tractable likelihoods for rigorous evaluation.
The state of the art has therefore shifted toward score-based diffusion models, which combine stable optimisation with likelihoods that are at least numerically tractable and come with explicit error bounds. Denoising Diffusion Probabilistic Models (DDPM) by Ho et al. (2020), together with the continuous-time score-SDE formulation of Song et al. (2021b), supply the theoretical foundation; in practice, likelihoods are obtained via discretised importance sampling rather than closed form, yet remain usable for model comparison. Finite-sample convergence guarantees—expressed as Wasserstein or KL error bounds under Lipschitz score error—have since been derived by Bortoli (2022) and Gao et al. (2025), providing theoretical guarantees that are currently unavailable for standard GAN objectives.
Building on these ideas, TimeGrad employs an autoregressive denoising architecture for multivariate prediction Rasul et al. (2021); ScoreGrad adapts continuous energy-based diffusion for improved sampling efficiency Yan et al. (2021); and CSDI applies conditional score matching to time-series imputation Tashiro et al. (2021). Subsequent advances such as TimeDiff Shen and Kwok (2023) and TSDiff Kollovieh et al. (2023) show that diffusion generators can outperform transformer baselines in forecasting, while FTS-Diffusion Huang et al. (2024a) focuses on high-fidelity unconditional synthesis and demonstrates downstream forecasting gains through data augmentation rather than direct head-to-head benchmarks. Compared with GANs, diffusion models thus offer tractable log-likelihoods, stronger convergence guarantees, and greater robustness, making them particularly well suited to the dynamic mean–variance problems addressed in the present work.
Generative models are moving beyond passive scenario engines and into the decision loop itself. Auxiliary-classifier GANs that condition on market states can learn return distributions whose samples feed directly into risk–return objectives, yielding portfolios that beat classical bootstrap or historical-window benchmarks Kim and Lee (2023). Diffusion-based methods advance the idea further: denoising-diffusion probabilistic models have been trained to generate realistic high-dimensional correlation matrices, enabling more faithful risk-parity and stress-testing exercises than traditional parametrics Kubiak et al. (2024), and score-based (continuous-time) diffusion model admits explicit statistical guarantee that addressing curse of dimensionality when there are hidden low-dimensional structure Chen et al. (2025).
Beyond using AI to generate synthetic scenarios, researchers have also turned to reinforcement learning (RL) for the allocation task itself. Modern RL systems learn portfolio policies end-to-end, treating rebalancing as a sequential control problem with risk-adjusted rewards. Empirical studies on cryptocurrencies and U.S. equities report multi-fold wealth growth and post-cost Sharpe ratios exceeding two Jiang et al. (2017), Cong et al. (2021). A recent survey consolidates these advances while highlighting ongoing challenges in sample efficiency, trading frictions, and risk-aware exploration Hambly et al. (2023).
Existing work either builds a market simulator or learns an RL allocation rule, but not both in one framework. We close this gap by first training a score-based diffusion model that produces state-conditional market scenarios, and then training a policy-gradient portfolio strategy on those simulated paths. A stability theorem ties the simulator’s adapted-Wasserstein error to the policy’s mean-variance performance, so improving the diffusion model keeps the RL performance near-optimal.
1.3 Organization of the paper
The rest of this paper is organized as follows. In Section 2 we present an adaptive training and sampling framework of score-based diffusion models, and provide a $\mathcal{A}\mathcal{W}_{2}$-quantification error bound between the output distribution and data distribution. In Section 3, we turn to a specific mean-variance portfolio selection problem, and establish the model stability in terms of the adapted Wasserstein metric (i.e., the value function is Lipschitz in the adapted Wasserstein metric). Section 4 focuses on applications of the theoretical results and, in particular, develops a policy gradient algorithm based on generative environment along with experimental results. 5 concludes the main context. Appendices include technical discussions and proofs of the main results in Section 2.
We will close this section by introducing frequently used notation in the paper.
1.4 Frequently used notation
In this work, the data is in the form of time-series with length $T\in\mathbb{N}$ and dimension $d\in\mathbb{N}$. We denote by $t\in\{1,2,\cdots,T\}$ the time index for the time-series. For general theoretical results, we use $X^{1:t}:=(X^{1},X^{2},\cdots,X^{t})$ ($Y^{1:t}:=(Y^{1},Y^{2},\cdots,Y^{t}$)) to denote random variables sampled from data distribution (approximated model, respectively), and $x^{1:t}$ ($y^{1:t}$) the realization. For portfolio selection problem, we use $S^{1:t}$ ($\tilde{S}^{1:t}$) and $s^{1:t}$ ($\tilde{s}^{1:t}$) instead. For the diffusion model, we use $\tau\in[0,\mathcal{T}]$ to denote the diffusive time, i.e., the time flow of the forward and reversed SDE, with $\mathcal{T}>0$ the maximal diffusive time.
For a probability $\mathbb{P}$ on the space of time-series data $\mathbb{R}^{dT}$, $\mathbb{P}_{1:t}$ is the joint distribution of $(X^{1},X^{2},\cdots X^{t})$, thus $\mathbb{P}_{1:T}=\mathbb{P}$. Moreover, $\mathbb{P}_{x^{1:t}}$ is the conditional distribution of $X^{t+1}$, conditioning on $X^{1:t}=x^{1:t}$. By convention, we use $\mathbb{P}$ to denote the original (real) model, hence the distribution of $X^{1:T}$ ($S^{1:T}$) and $\mathbb{Q}$ the alternative model, hence the distribution of $Y^{1:T}$ ($\tilde{S}^{1:T}$).
To present the quantification error bound of the score-based diffusion model, we use the adapted Wasserstein metric. More specifically, for two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathbb{R}^{dT}$, a coupling $\pi$ between $\mathbb{P}$ and $\mathbb{Q}$ is said to be causal in $x$ if for any $t\in\{1,2,\cdots,T\}$,
$\pi(Y^{t}\in\mathrm{d}y^{t}|X^{1:T}=x^{1:T})=\pi(Y^{t}\in\mathrm{d}y^{t}|X^{1:t}=x^{1:t}).$ (1.4)
A coupling $\pi$ is said to be bicausal if it is causal in both $x$ and $y$ (c.f. Backhoff et al. (2017), Bayraktar and Han (2025+)). Define $\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ to be the set of all bicausal couplings between $\mathbb{P}$ and $\mathbb{Q}$. With the choice of transport cost $|x^{1:T}-y^{1:T}|:=(\sum_{t=1}^{T}|x^{t}-y^{t}|^{2})^{1/2}$, we define the adapted Wasserstein metric as follows:
$\mathcal{A}\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q})=\bigg{(}\inf_{\pi\in\Pi_{\rm bc}}\int_{\mathbb{R}^{dT}\times\mathbb{R}^{dT}}|x^{1:T}-y^{1:T}|^{2}\pi(\mathrm{d}x^{1:T},\mathrm{d}y^{1:T})\bigg{)}^{1/2}.$ (1.5)
We use $K,C,c$ to denote generic constants which may vary from line to line.
2 Adaptive sampling via diffusion models
2.1 Conditional score matching
For simplicity, we take the OU process as the forward process of our diffusion model:
$\mathrm{d}X_{\tau}=-X_{\tau}\mathrm{d}\tau+\sqrt{2}\mathrm{d}B_{\tau}.$ (2.1)
The corresponding time-reversed SDE is
$\mathrm{d}\bar{X}_{\tau}=[\bar{X}_{\tau}+2\nabla\log p(\mathcal{T}-\tau,\bar{X}_{\tau})]\mathrm{d}t+\sqrt{2}\mathrm{d}\bar{B}_{\tau},$ (2.2)
where $p$ is the density function of the forward process (2.1) with certain initial conditions.
To facilitate our adaptive sampling method, we assume that we have access to a family of pre-trained score matching functions $s^{t}_{\theta}:[0,\mathcal{T}]\times\mathbb{R}^{dt}\to\mathbb{R}^{d}$, $t=1,2,\cdots,T$ satisfying the following conditional score matching error bounds:
Assumption 1.
For any $\tau\in(0,\mathcal{T}]$ and $t=1,2,\cdots,T-1$, we have
$\mathbb{E}_{X\sim p_{1}(\tau,\cdot)}|s^{1}_{\theta}(\tau,X)-\nabla_{x}\log p_{1}(\tau,X)|^{2}\leq\varepsilon^{2}_{\rm score},$ (2.3)
$\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\bigg{|}s^{t+1}_{\theta}(\tau,X^{1:t},X^{t+1}_{\tau})-\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|X^{1:t})\bigg{|}^{2}\leq\varepsilon^{2}_{\rm score}.$ (2.4)
Here, $p_{t+1}(\tau,\cdot|x^{1:t})$ is the probability density function of the forward process, with random initial condition $\mathbb{P}_{x^{1:t}}$.
We remark that (2.3) is the classical score error bound for the first marginal of $\mathbb{P}$; see Song et al. (2021b). (2.4) is crucial to our adaptive sampling scheme because it provides a proxy to $\nabla_{x}\log p_{t+1}(\tau,\cdot|X^{1:t})$, the score function conditioning on the realization $X^{1:t}$. However, (2.4) is not feasible directly, because we can not sample from the conditional probability, but rather, only have access to the joint distribution (data distribution). This issue can be resolved by the following denoising score-matching techniques, in which the error can be computed using only joint distribution. While the idea is borrowed from classical denoising score matching (e.g., Hyvärinen (2005)), we need to consider a conditional version and the proof is more complicated.
Proposition 2.1.
For any $t=1,2,\cdots,T-1$, the following two score-matching problems are equivalent:
$\min_{\theta}\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\bigg{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|X^{1:t})\bigg{|}^{2}$ (2.5)
$\min_{\theta}\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{X_{\tau}^{t+1}\sim\phi(\tau,\cdot|X^{t+1})}\bigg{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log\phi(\tau,X_{\tau}^{t+1}|X^{t+1})\bigg{|}^{2}.$ (2.6)
Here, $\phi(\tau,\cdot|x_{0})$ is the probability density function of the forward process, with initial condition $X_{0}=x_{0}$. In particular,
$\phi(\tau,x|x_{0})=(2\pi(1-e^{-2\tau}))^{-d/2}e^{-\frac{|x-x_{0}e^{-\tau}|^{2}}{2(1-e^{-2\tau})}}.$
Proof.
By definition, it is not hard to show that
$p_{t+1}(\tau,x|x^{1:t})=\int_{\mathbb{R}^{d}}\phi(\tau,x|x_{0})\mathbb{P}_{x^{1:t}}(\mathrm{d}x_{0}).$ (2.7)
The term to be minimized in (2.5) can thus be calculated by
$\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\bigg{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|X^{1:t})\bigg{|}^{2}$ (2.8)
$=$ $\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\big{|}^{2}$ (2.9)
$+\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\big{|}\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|X^{1:t})\big{|}^{2}$ (2.10)
$-2\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\big{[}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\cdot\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|X^{1:t})\big{]}$ (2.11)
$=$ $\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\big{|}^{2}+C$ (2.12)
$-2\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\int_{\mathbb{R}^{d}}s^{t+1}_{\theta}(\tau,X^{1:t},x)\cdot\nabla_{x}p_{t+1}(\tau,x|X^{1:t})\mathrm{d}x$ (2.13)
$=$ $\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\big{|}^{2}+C$ (2.14)
$-2\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\int_{\mathbb{R}^{d}}s^{t+1}_{\theta}(\tau,X^{1:t},x)\cdot\nabla_{x}\int_{\mathbb{R}^{d}}\phi(\tau,x|x_{0})\mathbb{P}_{X^{1:t}}(\mathrm{d}x_{0})\mathrm{d}x$ (2.15)
$=$ $\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\big{|}^{2}+C$ (2.16)
$-2\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X^{t+1}\sim\mathbb{P}_{X^{1:t}}}\mathbb{E}_{X_{\tau}^{t+1}\sim\phi(\tau,\cdot|X^{t+1})}\big{[}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\cdot\nabla_{x}\phi(\tau,X_{\tau}^{t+1}|X^{t+1})\big{]}$ (2.17)
$=$ $\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|X^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\big{|}^{2}+C$ (2.18)
$-2\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{X_{\tau}^{t+1}\sim\phi(\tau,\cdot|X^{t+1})}\big{[}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\cdot\nabla_{x}\phi(\tau,X_{\tau}^{t+1}|X^{t+1})\big{]}$ (2.19)
$=$ $\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{X_{\tau}^{t+1}\sim\phi(\tau,\cdot|X^{t+1})}\bigg{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log\phi(\tau,X_{\tau}^{t+1}|X^{t+1})\bigg{|}^{2}+C.$ (2.20)
Here, $C$ is a generic constant independent from $\theta$. The proof is completed because (2.5) and (2.6) attain the same minimizer. ∎
Remark 1.
Because the explicit form of $\phi(\tau,x|x_{0})$, $X^{t+1}_{\tau}\sim\phi(\tau,\cdot|X^{t+1})$ is equivalent to $X^{t+1}_{\tau}=X^{t+1}e^{-\tau}+\sqrt{1-e^{-2\tau}}\mathbf{z}$, where $\mathbf{z}\sim\mathcal{N}(0,I)$. Therefore, (2.6) is further equivalent to
$\min_{\theta}\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{\mathbf{z}\sim\mathcal{N}(0,I)}\big{|}\sqrt{1-e^{-2\tau}}s^{t+1}_{\theta}\big{(}\tau,X^{1:t},X^{t+1}e^{-\tau}+\sqrt{1-e^{-2\tau}}\mathbf{z}\big{)}+\mathbf{z}\big{|}^{2}.$ (2.21)
In the implementation, $s_{\theta}$ is trained so as to minimize the empirical version of (2.21), with an additional integration in $\tau$ to account for different diffusive time; see (4.1).
2.2 An adaptive sampling method
Suppose the real model of time-series data is $\mathbb{P}$, and we have a pre-trained diffusion model $s_{\theta}^{t}$ satisfies (2.3) and (2.4). We note that, fixing $x^{1:t}$, $s^{t+1}_{\theta}(\tau,x^{1:t},\cdot)$, serves as an approximation of the conditional score function $\nabla_{x}\log p(\tau,\cdot|x^{1:t})$. This motives us to sample from the output distribution adaptively; see Algorithm 1.
Algorithm 1 Adaptive sampling
Inputs: pre-trained approximated score functions $s^{t}_{\theta}$, $t=1,2,\cdots,T$; backward SDE simulators $\{Y_{s}\}_{0\leq s\leq\mathcal{T}}$.
Initialization: samples from the noise: $\{(Z_{(n)}^{1},Z_{(n)}^{2},\cdots,Z_{(n)}^{T})\}_{n=1}^{N}\sim\mathcal{N}({\bf 0}_{\mathbb{R}^{dT}},I_{dT\times dT})$.
1:for $n=1$ to NN do
2:   With initial condition $\bar{Y}_{0}=Z_{(n)}^{1}$, run the backward SDE with score matching function $s^{\theta}_{1}$. , i.e.,
$\mathrm{d}\bar{Y}_{\tau}=[\bar{Y}_{\tau}+2s_{\theta}^{1}(\mathcal{T}-\tau,\bar{Y}_{\tau})]\mathrm{d}\tau+\sqrt{2}\bar{B}_{\tau}.$ (2.22)
3:   $y_{(n)}^{1}\leftarrow\bar{Y}_{\mathcal{T}}$.
4:  for $t=1$ to T1T-1 do
5:   $y^{1:t}\leftarrow y^{1:t}_{(n)}$
6:    With initial condition $Y_{0}=Z_{(n)}^{t+1}$, run the backward SDE with score matching function $(\tau,x)\mapsto s^{t+1}_{\theta}(\tau,y^{1:t},x)$. , i.e,
$\mathrm{d}\bar{Y}_{\tau}=[\bar{Y}_{\tau}+2s_{\theta}^{t+1}(\mathcal{T}-\tau,y^{1:t},\bar{Y}_{\tau})]\mathrm{d}\tau+\sqrt{2}\bar{B}_{\tau}.$ (2.23)
7:    $y_{(n)}^{t+1}\leftarrow\bar{Y}_{\mathcal{T}}$.
8:  end for
9:end for
Output: $\{(y_{(n)}^{1},y_{(n)}^{2},\cdots,y_{(n)}^{T})\}_{n=1}^{N}$.
We now proceed to prove that the sampling method described by Algorithm 1 gives a distribution $\mathbb{Q}$ close to $\mathbb{P}$ in $\mathcal{A}\mathcal{W}_{2}$. Our proof relies on the existing total variation bounds of diffusion models between conditional distributions and an uniform tail estimate of SDEs. We need the following technical assumptions for our results.
Assumption 2 (Assumptions on the data distribution).
  1. 1.
    There exists a constant $L>0$ such that for any $t\in\{1,2,\cdots,T\}$, $\tau\in[0,\mathcal{T}]$ and $x^{1:t},y^{1:t}\in\mathbb{R}^{dt}$,
    $|\nabla_{x}\log p_{1}(\tau,x)-\nabla_{x}\log p_{1}(\tau,y)|\leq L|x-y|,$ (2.24)
    $|\nabla_{x}\log p_{t+1}(\tau,x|x^{1:t})-\nabla_{x}\log p_{t+1}(\tau,y|y^{1:t})|\leq L(|x^{1:t}-y^{1:t}|+|x-y|).$ (2.25)
  2. 2.
    For some $c>0$, $\mathbb{E}_{\mathbb{P}_{1}}e^{c|X_{1}|}<\infty$, and for $t=\{1,2,\cdots,T-1\}$,
    $\sup_{x^{1:t}\in\mathbb{R}^{dt}}\mathbb{E}_{\mathbb{P}_{x^{1:t}}}e^{c|X^{t+1}|}<\infty.$ (2.26)
Assumption 3 (Assumptions on the approximating network).
  1. 1.
    There exists a constant $L>0$ such that for any $t\in\{1,2,\cdots,T\}$, $\tau\in[0,\mathcal{T}]$ and $x^{1:t},y^{1:t}\in\mathbb{R}^{dt}$,
    $s^{1}_{\theta}(\tau,x)-s^{1}_{\theta}(\tau,y)\leq L|x-y|,$ (2.27)
    $s^{t+1}_{\theta}(\tau,x^{1:t},x)-s^{t+1}_{\theta}(\tau,y^{1:t},y)\leq L(|x^{1:t}-y^{1:t}|+|x-y|).$ (2.28)
  2. 2.
    There exist constants $M_{\rm disp},\delta>0$ such that for any for any $t\in\{1,2,\cdots,T\}$, $\tau\in[0,\mathcal{T}]$, $x^{1:t}\in\mathbb{R}^{dt}$ and $x\in\mathbb{R}^{d}$,
    $2x\cdot s^{1}_{\theta}(\tau,x)\leq-(1+\delta)|x|^{2}+{M_{\rm disp}},$ (2.29)
    $2x\cdot s^{t+1}_{\theta}(\tau,x^{1:t},x)\leq-(1+\delta)|x|^{2}+{M_{\rm disp}}.$ (2.30)
Typically, it is advisable to refrain from making assumptions about the approximation network, as such assumptions may potentially compromise its approximating power. However, in this paper, we demonstrate that by adhering to Assumption 2, which are all assumptions pertaining to the data distribution, we can select the approximating network $s_{\theta}$ such that it satisfies Assumption 3 (with appropriately chosen $M_{\rm disp}$), while maintaining the approximation errors (2.3) and (2.4). See Proposition 2.2 below. To avoid deviating from the primary objective of this paper, we present detailed discussions on this result in Appendix A.
Proposition 2.2.
Suppose Assumption 2 holds. For any $\varepsilon_{\rm score}>0$, there exists a $s_{\theta}$ satisfying Assumption 1 as well as Assumption 3 with $M_{\rm disp}\sim\log(1/\varepsilon_{\rm score})$.
We are now ready to present one of our main results: the quantitative error bounds between output distribution of the adaptive sampling method, and the real data distribution.
Theorem 2.3.
For $t\geq 1$, denote by $\{X^{x^{1:t}}_{\tau}\}_{\tau\in[0,\mathcal{T}]}$ the forward process with initial condition $X^{x^{1:t}}_{0}\sim\mathbb{P}_{x^{1:t}}$, and $\{Y^{y^{1:t}}_{\tau}\}_{\tau\in[0,\mathcal{T}]}$ the backward process with initial condition $Y^{y^{1:t}}_{0}\sim\mathcal{N}(0_{\mathbb{R}^{d}},I_{d\times d})$ and score matching function $(\tau,x)\mapsto s^{t}_{\theta}(\tau,y^{1:t},x)$, i.e., $Y^{y^{1:t}}$ satisfies (2.23). Define $\mathbb{Q}^{\mathcal{T}}_{y^{1:t}}=\mathcal{L}(Y^{y^{1:t}}_{\mathcal{T}})$. Moreover, define $\mathbb{Q}^{\mathcal{T}}_{1}=\mathcal{L}(Y_{\mathcal{T}})$ with $Y$ satisfying (2.22).
  • (1)
    There exists a constant $c>0$ such that, for $t=1,2,\cdots,T-1$ we have
    $\mathcal{W}_{2}^{2}\big{(}\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}}\big{)}\leq$ $C\Big{(}\alpha(\mathcal{T})+\mathcal{T}^{2}\mathcal{E}(x^{1:t})^{1/2}+\mathcal{T}^{5/2}|x^{1:t}-y^{1:t}|\Big{)},$ (2.31)
    where
    $\mathcal{E}(x^{1:t}):=\int_{0}^{\mathcal{T}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|x^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,x^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|x^{1:t})\big{|}^{2}\mathrm{d}\tau,$ (2.32)
    $\alpha(\mathcal{T}):=\mathcal{T}^{2}e^{-\mathcal{T}}+e^{-c\mathcal{T}+cM_{\rm disp}}.$ (2.33)
  • (2)
    Define $\mathbb{Q}^{\mathcal{T}}$ to be the joint distribution with regular conditional kernels $\{\mathbb{Q}^{\mathcal{T}}_{x^{1:t}}:x^{1:t}\in\mathbb{R}^{dt},t=1,2,\cdots,T-1\}$, i.e.,
    $\mathbb{Q}^{\mathcal{T}}(\mathrm{d}x^{1},\mathrm{d}x^{2},\cdots,\mathrm{d}x^{T}):=\mathbb{Q}^{\mathcal{T}}_{1}(\mathrm{d}x^{1})\mathbb{Q}^{\mathcal{T}}_{x^{1}}(\mathrm{d}x^{2})\cdots\mathbb{Q}^{\mathcal{T}}_{x^{1:t}}(\mathrm{d}x^{t+1})\cdots\mathbb{Q}^{\mathcal{T}}_{x^{1:T-1}}(\mathrm{d}x^{T}).$ (2.34)
    Then, if $\mathcal{T}>1$, $\alpha(\mathcal{T})<1$ and $\varepsilon_{\rm score}<1$, we have
    $\mathcal{A}\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq C\Big{(}\mathcal{T}^{\frac{5T}{2}}\varepsilon_{\rm score}^{1/2^{T-1}}+\mathcal{T}^{\frac{5(T-1)}{2}}\alpha(\mathcal{T})^{1/2^{T-1}}\Big{)}.$ (2.35)
Proof.
See Appendix B. ∎
Several remarks on Theorem 2.3 follows.
Remark 2.
  • (1)
    By Proposition 2.2, with vanishing $\varepsilon_{\rm score}$ we can take $M_{\rm disp}\sim\log(1/\varepsilon_{\rm score})$, thus $\alpha(\mathcal{T})\leq\mathcal{T}^{2}e^{-c\mathcal{T}+cM_{\rm disp}}\leq\mathcal{T}^{2}\varepsilon_{\rm score}^{-1}e^{-c\mathcal{T}}$. In this case (2.35) reads
    $\mathcal{A}\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq C\Big{(}\mathcal{T}^{\frac{5T}{2}}\varepsilon_{\rm score}^{1/2^{T-1}}+\mathcal{T}^{\frac{5(T-1)}{2}+\frac{1}{2^{T-2}}}e^{-c\mathcal{T}/2^{T-1}}\varepsilon_{\rm score}^{-1/2^{T-1}}\Big{)}.$ (2.36)
    Thus, taking $\varepsilon_{\rm score}=\mathcal{T}^{-52^{T-3}+1}e^{-c\mathcal{T}/2}$, (2.36) becomes
    $\mathcal{A}\mathcal{W}^{2}_{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq C\mathcal{T}^{\frac{5T}{2}-\frac{5}{4}+\frac{1}{2^{T-1}}}e^{-\frac{c\mathcal{T}}{2^{T}}}.$ (2.37)
    which provides a convergence guarantee when $\mathcal{T}\to\infty$ (equivalently, $\varepsilon_{\rm score}\to 0$). Even in the static data scenario, achieving convergence under the Wasserstein metric for the diffusion model is recognized as challenging; see Remark 1 of Kwon et al. (2022). Under various technical assumptions on the data distribution, such as strong log-concavity, there are Wasserstein convergence results available (e.g., Gao et al. (2025), Tang and Zhao (2024)), and it is possible to bound $\mathcal{A}\mathcal{W}_{2}$ by $\mathcal{W}_{2}$ up to some explicit error terms; see Blanchet et al. (2024). However, with this conventional approach, obtaining bounds for conditional distributions requires imposing strong log-concavity on every conditional distribution of $\mathbb{P}$. We circumvent these assumptions and replace them with certain disspative conditions of the networks, which are attained by carefully designing the network structure (i.e., Assumptions 3).
  • (2)
    Taking $T=1$ in (2.31) or (2.35), we obtain from (2.36) a Wasserstein bound for static data generation:
    $\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq C(\mathcal{T}^{\frac{5}{2}}\varepsilon_{\rm score}+\mathcal{T}^{2}e^{-c\mathcal{T}}\varepsilon_{\rm score}^{-1}),$
    which is interesting in its own right. This is because when $T=1$, Assumption 2 is reduced to the usual Lipschitz continuity of score functions, and, again, we do not rely on any structural assumptions on $\mathbb{P}$, such as strongly log concavity. This is achieved by bounding Wasserstein metric by total variation, up to some tail error (Blanchet et al. (2024), Proposition 20, and Villani (2008), Theorem 6.15), and making use of the uniform tail estimates of SDEs with dissipative coefficients. The price to pay is that the noise approximation error term $\mathcal{T}^{2}e^{-c\mathcal{T}}\varepsilon_{\rm score}^{-1}$ explodes when $\varepsilon_{\rm score}\to 0$.
  • (3)
    It is evident that, as outlined in Algorithm 1, $\{y_{(n)}^{1}\}_{n=1}^{N}$ is sampled from the distribution $\mathbb{Q}_{1}^{\mathcal{T}}$. Furthermore, given the sequence $y_{(n)}^{1:t}$, the next sample $y_{(n)}^{t+1}$ is drawn from the distribution $\mathbb{Q}^{\mathcal{T}}_{y_{(n)}^{1:t}}$. By definition, $\{y_{(n)}^{1},y_{(n)}^{2},\cdots,y_{(n)}^{T}\}_{n=1}^{N}$ is one sample from the distribution $\mathbb{Q}^{\mathcal{T}}$, thereby establishing that $\mathbb{Q}^{\mathcal{T}}$ is the output distribution of Algorithm 1. Consequently, Theorem 2.3 provides $\mathcal{A}\mathcal{W}_{2}$ bounds between the output distribution of Algorithm 1 and the data distribution.
  • (4)
    Due to the construction of $\mathbb{Q}^{\mathcal{T}}$ (as defined in (2.34)), for a random variable $Y^{1:T}\sim\mathbb{Q}^{\mathcal{T}}$, $\mathbb{Q}^{\mathcal{T}}_{y^{1:t}}$ is the conditional distribution of $Y^{t+1}$ given $Y^{1:t}=y^{1:t}$. Consequently, it can be sampled by selecting the score matching function $s^{t+1}_{\theta}$. In essence, Algorithm 1 can also be utilized for conditional sampling, provided that $y^{1:t}$ is included as input, and the sampling process starts with $Z_{(n)}^{t+1}$.
A useful corollary of Theorem 2.3 is that we can also bound the distance between conditional distributions.
Corollary 2.4.
For $t=1,2,\cdots,T-1$ we have
$\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}[\mathcal{W}^{2}_{2}(\mathbb{P}_{X^{1:t}},\mathbb{Q}_{X^{1:t}})]\leq C(\mathcal{T}^{\frac{5}{2}}\varepsilon_{\rm score}+\alpha(\mathcal{T})),$ (2.38)
where $\alpha(\mathcal{T})$ is given by (2.33).
Consequently, for any $\varepsilon,p>0$, if $\mathcal{T}$ is sufficiently large such that $\alpha(\mathcal{T})<\varepsilon p/(2C)$, and $\varepsilon_{\rm score}<\frac{\varepsilon p\mathcal{T}^{-\frac{5}{2}}}{2C}$, then with probability at least 1p1-p, we have
$\mathcal{W}_{2}^{2}(\mathbb{P}_{X^{1:t}},\mathbb{Q}_{X^{1:t}})\leq\varepsilon,$ (2.39)
provided that the prior $X^{1:t}$ is randomly sampled from $\mathbb{P}_{1:t}$.
3 Stability of mean-variance portfolio selection problems
In this section, we first formulate the dynamic mean-variance portfolio selection problem, allowing general constraint set $K$. We employ a dual argument and then establish the dynamic programming principle for the dual quadratic hedging problem in Subsection 3.2. In Subsection 3.3, in preparation for the application of our diffusion model, we provide a model stability result with $\mathcal{A}\mathcal{W}_{2}$ metric.
Suppose the stock prices $(S^{1},S^{2},\cdots,S^{T})$ satisfies some model $\mathbb{P}$. Suppose the risk-free interest rate is 0. For a investment strategy $\vartheta$ (in some admissible or constraint set), the value associated with it is $(\vartheta\cdot S)_{t}:=\sum_{l=1}^{t-1}\vartheta_{l}^{\mathbf{t}}(S^{l+1}-S^{l})$. We want to maximize the mean-variance objective:
$v(\mathbb{P};\vartheta):=\mathbb{E}_{\mathbb{P}}[(\vartheta\cdot S)_{T}]-\frac{\gamma}{2}\mathrm{Var}_{\mathbb{P}}[(\vartheta\cdot S)_{T}].$ (3.1)
$v^{*}(\mathbb{P}):=\sup_{\vartheta}v(\mathbb{P};\vartheta).$ (3.2)
Using a classical result from literature, if we know the real model $\mathbb{P}$, it is sufficient to solve the quadratic hedging problem:
$V(\mathbb{P},c)=\min_{\vartheta}\mathbb{E}_{\mathbb{P}}\big{[}|c-(\vartheta\cdot S)_{T}|^{2}\big{]}.$ (3.3)
for some $c>0$. This is because, for $\vartheta$ such that $\mathbb{E}_{\mathbb{P}}[(\vartheta\cdot S)_{T}]>0$,
$v(\mathbb{P};\vartheta)=\sup_{a\in\mathbb{R}_{+}}\bigg{\{}-\frac{\gamma}{2}\mathbb{E}_{\mathbb{P}}\Big{[}\Big{|}(\vartheta\cdot S)_{T}-\Big{(}\frac{1}{\gamma}+a\Big{)}\Big{|}^{2}\Big{]}+\frac{1}{2\gamma}+a\bigg{\}}.$ (3.4)
Consequently, we have the duality relation between these two value functions:
$v^{*}(\mathbb{P})=\sup_{a\in\mathbb{R}_{+}}\bigg{\{}-\frac{\gamma}{2}V\bigg{(}\mathbb{P},\frac{1}{\gamma}+a\bigg{)}+\frac{1}{2\gamma}+a\bigg{\}}.$ (3.5)
Remark 3.
If we consider cone constraint on portfolio, then it is sufficient to study (3.3) with $c=1$, and there is an explicit relation between solutions of these two problems. For generality, we choose to consider a generic convex constraint set $K$. In particular, we can consider $K$ to be bounded.
3.1 Technical preparations
This subsection presents several technical results regarding the quadratic hedging problem 3.3. There is a vast set of works concerning this type of problem, see Labbé and Heunis (2007), Czichowsky and Schweizer (2012), Černý et al. (2024), among others. Here, we consider a general bounded convex constraint set, making some classical results not directly applicable. For simplicity, we also make stronger assumptions on the market model and admissible portfolios. Hence, we provide proofs for completeness.
Denote $\triangle S^{t}=S^{t+1}-S^{t}$. Suppose $\mathbb{F}$ is the natural filtration of $S$, and $\mathcal{F}_{1}$ is trivial.
Assumption 4.
$\mathbb{E}_{\mathbb{P}}\big{[}|S^{t+1}|^{2}\big{|}\mathcal{F}_{t}\big{]}$ is uniformly bounded for $t=1,2,\cdots T-1$. Moreover, $\mathbb{P}(S^{1}=s^{1})=1$ for some positive constant $s^{1}$.
Assumption 5.
There exists $\delta\in(0,1)$ such that,
$\bigg{(}\mathbb{E}_{\mathbb{P}}\big{[}\triangle S^{t}|\mathcal{F}_{t}\big{]}\bigg{)}\bigg{(}\mathbb{E}_{\mathbb{P}}\big{[}\triangle S^{t}|\mathcal{F}_{t}\big{]}\bigg{)}^{\mathbf{t}}\leq(1-\delta)\mathbb{E}_{\mathbb{P}}\big{[}\triangle S^{t}(\triangle S^{t})^{\mathbf{t}}|\mathcal{F}_{t}\big{]},\forall t=1,2,\cdots,T-1.$ (3.6)
In other words, for any $\phi\in\mathbb{R}^{d}$,
$\Big{|}\big{(}\mathbb{E}_{\mathbb{P}}\big{[}\phi^{\mathbf{t}}\triangle S^{t}|\mathcal{F}_{t}\big{]}\big{)}\Big{|}^{2}\leq(1-\delta)\mathbb{E}_{\mathbb{P}}\big{[}|\phi^{\mathbf{t}}\triangle S^{t}|^{2}|\mathcal{F}_{t}\big{]}$ (3.7)
Definition 3.1.
$L^{2}_{\mathbb{P}}(S)$ is the set of $\mathbb{F}$-adapted, $\mathbb{R}^{d}$-valued processes $\{\vartheta_{t}\}_{t=1}^{T-1}$ such that
$\mathbb{E}_{\mathbb{P}}\big{[}|\vartheta^{\mathbf{t}}_{t}\triangle S^{t}|^{2}\big{]}<\infty,\forall t=1,2,\cdots,T-1.$ (3.8)
Remark 4.
By Assumption 4, constant strategies are in $L^{2}_{\mathbb{P}}(S)$.
Remark 5.
From Schweizer (1995) we know (with apparent extension to multi-dimensional case), (3.6) ensures that the set of unconstrained admissible terminal wealth $G(L^{2}(S)):=\{(\vartheta\cdot S)_{T}:\vartheta\in L^{2}(S)\}$ is a closed subspace in $L^{2}(\mathbb{P})$. Consequently, the optimizer of unconstrained problem (3.3) exists. Moreover, by strong convexity of quadratic function, the optimizer is unique in value, i.e., $(\vartheta\cdot S)_{T}=(\tilde{\vartheta}\cdot S)_{T}$, $\mathbb{P}$-a.s., for any two optimizers $\vartheta$ and $\tilde{\vartheta}$.
In our work, we want to consider portfolio constraint.
Definition 3.2.
For a closed convex subset $K\subset\mathbb{R}^{d}$, define $\Theta_{K,\mathbb{P}}=\{\vartheta\in L^{2}(S):\vartheta_{t}\in K,\ \mathbb{P}{\rm-a.s.}\}$, and $G(\Theta_{K,\mathbb{P}}):=\{(\vartheta\cdot S)_{T}:\vartheta\in\Theta_{K,\mathbb{P}}\}$. In this paper, we assume that $K$ is bounded.
Lemma 3.3.
Denote $\Pi^{S}_{t}:=\sqrt{\mathbb{E}_{\mathbb{P}}[\triangle S^{t}(\triangle S^{t})^{\mathbf{t}}|\mathcal{F}_{t}]},t=1,2,\cdots,T-1$. Then for any $\vartheta,\bar{\vartheta}\in L^{2}(S)$, $(\vartheta\cdot S)_{t}=(\bar{\vartheta}\cdot S)_{t}$ for any $t=1,2,\cdots,T-1$ if and only if $\Pi^{S}_{t}\vartheta=\Pi^{S}_{t}\bar{\vartheta}$ for any $t=1,2,\cdots,T-1$.
Proof.
From definition we know $(\vartheta\cdot S)_{t}=(\bar{\vartheta}\cdot S)_{t}$ for any $t=1,2,\cdots,T-1$ if and only if $\vartheta^{\mathbf{t}}\triangle S^{t}=\bar{\vartheta}^{\mathbf{t}}\triangle S^{t}$, which in turn is equivalent to
$\mathbb{E}_{\mathbb{P}}[|(\vartheta-\bar{\vartheta})_{t}^{\mathbf{t}}\triangle S^{t}|^{2}]=0.$ (3.9)
The proof is completed by observing that
$0=$ $\mathbb{E}_{\mathbb{P}}[|(\vartheta_{t}-\bar{\vartheta}_{t})^{\mathbf{t}}\triangle S^{t}|^{2}]$ (3.10)
$=$ $\mathbb{E}_{\mathbb{P}}\big{[}(\vartheta_{t}-\bar{\vartheta}_{t})^{\mathbf{t}}\triangle S^{t}(\triangle S^{t})^{\mathbf{t}}(\vartheta_{t}-\bar{\vartheta}_{t})\big{]}$ (3.11)
$=$ $\mathbb{E}_{\mathbb{P}}\big{[}(\vartheta_{t}-\bar{\vartheta}_{t})^{\mathbf{t}}\mathbb{E}_{\mathbb{P}}[\triangle S^{t}(\triangle S^{t})^{\mathbf{t}}|\mathcal{F}_{t}](\vartheta_{t}-\bar{\vartheta}_{t})\big{]}$ (3.12)
$=$ $\mathbb{E}_{\mathbb{P}}\big{[}(\vartheta_{t}-\bar{\vartheta}_{t})^{\mathbf{t}}\Pi^{S}_{t}(\Pi^{S}_{t})^{\mathbf{t}}(\vartheta_{t}-\bar{\vartheta}_{t})\big{]}$ (3.13)
$=$ $\mathbb{E}_{\mathbb{P}}[|\Pi^{S}_{t}(\vartheta_{t}-\bar{\vartheta}_{t})|^{2}].$ (3.14)
Proposition 3.4.
$G(\Theta_{K,\mathbb{P}})$ is a closed convex subset of $L^{2}(\mathbb{P})$.
Proof.
We only need to prove the closedness. Suppose $(\vartheta^{n}\cdot S)_{T}\to(\vartheta\cdot S)_{T}$ in $L^{2}(\mathbb{P})$ for some $\vartheta^{n}\in\Theta_{K,\mathbb{P}}$, and $\vartheta\in L^{2}(S)$. Consider the semimartingale decomposition of $S$, $S^{t}=s^{1}+A_{t}+M_{t}$, where $\triangle M_{t}:=S^{t+1}-\mathbb{E}_{\mathbb{P}}[S^{t+1}|\mathcal{F}_{t}]$, $\triangle A_{t}:=\mathbb{E}_{\mathbb{P}}[\triangle S^{t}|\mathcal{F}_{t}]$, $t=1,2,\cdots,T-1$. For any $t=1,2,\cdots,T-1$, we have
$(\vartheta\cdot S)_{t+1}=(\vartheta\cdot S)_{t}+\vartheta_{t}^{\mathbf{t}}\triangle A_{t}+\vartheta_{t}^{\mathbf{t}}\triangle M_{t}.$ (3.15)
Because the third term is orthogonal to the first two, we have
$\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}\big{(}(\vartheta^{n}-\vartheta)\cdot S\big{)}_{t+1}\big{|}^{2}\Big{]}=\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}((\vartheta^{n}-\vartheta)\cdot S)_{t}+(\vartheta^{n}_{t}-\vartheta_{t})^{\mathbf{t}}\triangle A_{t}\big{|}^{2}\Big{]}+\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}(\vartheta^{n}_{t}-\vartheta_{t})^{\mathbf{t}}\triangle M_{t}\big{|}^{2}\Big{]}.$ (3.16)
Starting from t=T1t=T-1, because $(\vartheta^{n}\cdot S)_{T}\to(\vartheta\cdot S)_{T}$ in $L^{2}(\mathbb{P})$, we know $\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\triangle M_{T-1}\big{|}^{2}\Big{]}\to 0$. From (3.6), it is not hard to derive
$\mathbb{E}_{\mathbb{P}}[|\Pi^{S}_{T-1}(\vartheta^{n}_{T-1}-\vartheta_{T-1})|^{2}]=$ $\mathbb{E}_{\mathbb{P}}\Big{[}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\mathbb{E}_{\mathbb{P}}[\triangle S^{T-1}(\triangle S^{T-1})^{\mathbf{t}}|\mathcal{F}_{T-1}](\vartheta^{n}_{T-1}-\vartheta_{T-1})\Big{]}$ (3.17)
$\leq\frac{\delta}{1-\delta}\mathbb{E}_{\mathbb{P}}\Big{[}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\mathbb{E}_{\mathbb{P}}[\triangle M_{T-1}\triangle M_{T-1}^{\mathbf{t}}|\mathcal{F}_{T-1}](\vartheta^{n}_{T-1}-\vartheta_{T-1})\Big{]}$ (3.18)
$=\frac{\delta}{1-\delta}\mathbb{E}_{\mathbb{P}}\Big{[}\big{|}(\vartheta^{n}_{T-1}-\vartheta_{T-1})^{\mathbf{t}}\triangle M_{T-1}\big{|}^{2}\Big{]}$ (3.19)
$\to 0.$ (3.20)
Thus $\Pi^{S}_{T-1}\vartheta^{n}_{T-1}\to\Pi^{S}_{T-1}\vartheta_{T-1}$ in $L^{2}(\mathbb{P})$. Equivalently, we have $(\vartheta^{n}_{T-1})^{\mathbf{t}}\triangle S^{T-1}\to\vartheta_{T-1}^{\mathbf{t}}\triangle S^{T-1}$ in $L^{2}(\mathbb{P})$. By using (3.15) and repeating the arguments, we actually conclude that $\Pi^{S}_{t}\vartheta^{n}_{t}\to\Pi^{S}_{t}\vartheta_{t}$ in $L^{2}(\mathbb{P})$ for any $t=1,2,\cdots,T-1$, hence up to a subsequence, the convergence is almost surely. Because $K$ is closed convex, we can define the projection operator $\Pi_{K}:\mathbb{R}^{d}\to K$. Clearly $\Pi^{S}_{t}\Pi_{K}(\mathbb{R}^{d})$ is a closed subset of $\mathbb{R}^{d}$, $\mathbb{P}$-almost surely. Therefore, there exists a $\vartheta^{\prime}$ such that $\Pi^{S}_{t}\Pi_{K}\vartheta^{\prime}_{t}=\Pi^{S}_{t}\vartheta_{t}$. By Lemma 3.3, $(\vartheta\cdot S)=(\Pi_{K}\vartheta^{\prime}\cdot S)$, thus $(\vartheta\cdot S)_{T}\in G(\Theta_{K,\mathbb{P}})$. ∎
Corollary 3.5.
There exists an optimizer in $\Theta(K)$ to the quadratic hedging problem (3.3), and the optimizer is unique in value, i.e., $(\vartheta\cdot S)_{T}=(\tilde{\vartheta}\cdot S)_{T}$, $\mathbb{P}$-a.s., for any two optimizers $\vartheta$ and $\tilde{\vartheta}$.
3.2 Dynamic programming
To derive DPP, we define a conditional version of (3.3) (for $t=1,2,\cdots,T$):
$V(t,w,s^{1:t}):=\inf_{\vartheta\in\Theta^{t,s^{1:t}}_{K,\mathbb{P}}}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\bigg{[}\bigg{|}c-w-\sum_{l=t}^{T-1}\vartheta_{l}^{\mathbf{t}}\triangle S^{l}\bigg{|}^{2}\bigg{]}.$ (3.21)
Definition 3.6.
$\Theta^{t,s^{1:t}}_{K,\mathbb{P}}$ is the set of processes $\vartheta_{l}$, $l=t,\cdots,T-1$, which are valued in $K$, $\mathbb{P}_{s^{1:t}}$-a.s., and are adapted to $\{\mathcal{F}^{t}_{l}\}_{l=t}^{T-1}$, with $\mathcal{F}^{t}_{l}=\sigma(S^{t+1},\cdots,S^{l})$ [^1] By convention we assume $\mathcal{F}^{t}_{t}$ is trivial..
Then, assuming $(\vartheta\cdot S)_{1}=0$ and $S^{1}=s^{1}$ is a constant, we have
$\inf_{\vartheta\in\Theta_{K,\mathbb{P}}}\mathbb{E}_{\mathbb{P}}\big{[}|c-(\vartheta\cdot S)_{T}|^{2}\big{]}=V(1,0,s^{1}).$ (3.22)
Note that, (3.6) is written in terms of conditional probability, in which $\mathcal{F}_{t}=\mathcal{F}^{0}_{t}=\sigma(S^{1},\cdots,S^{t})$. It is thus clear from Subsection 3.1 that (3.21) also has an optimizer that is unique in value.
Theorem 3.7.
For (3.21) we have the following DPP:
$V(T,w,s^{1:T})=|c-w|^{2},$ (3.23)
$V(t,w,s^{1:t})=\inf_{\theta\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}t+1,w+\theta^{\mathbf{t}}\triangle S^{t},(s^{1:t},S^{t+1})\Big{)}\Big{]},$ (3.24)
and for the initial problem $V(1,0,s^{1})$, an optimizer is given recursively by $\hat{\vartheta}_{t}=\hat{\theta}(t,(\hat{\vartheta}\cdot S)_{t},S^{1:t})$, where
$\hat{\theta}(t,w,s^{1:t})=\operatorname*{arg\,min}_{\theta\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}t+1,w+\theta^{\mathbf{t}}\triangle S^{t},(s^{1:t},S^{t+1})\Big{)}\Big{]}.$ (3.25)
Proof.
(3.23) is clear. Because the optimizer of (3.21) exists, we pick one such optimizer $\hat{\vartheta}_{l}\in\Theta^{t,s^{1:t}}_{K,\mathbb{P}}$, $l=t,t+1,\cdots,T-1$. By the definition of $\mathcal{F}^{t}_{l}$, $\hat{\vartheta}_{l}$ can be written as deterministic functions $\theta(l,S^{t+1},\cdots,S^{l})$. For any $s^{t+1}$, the random variable $\tilde{\vartheta}^{s^{t+1}}_{l}$ given by $\omega\mapsto\theta(l,s^{t+1},S^{t+2},\cdots,S^{l})$ is $\mathcal{F}^{t+1}_{l}$ measurable. Thus, $\{\tilde{\vartheta}_{l}^{s^{t+1}}\}_{l=t+1}^{T-1}\in\Theta^{t+1,s^{1:t+1}}_{K,\mathbb{P}}$. Consequently,
$V(t,w,s^{1:t})$ $=\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\mathbb{E}_{\mathbb{P}_{s^{1:t},S^{t+1}}}\bigg{[}\bigg{|}1-w-\hat{\vartheta}^{\mathbf{t}}_{t}\triangle S^{t}-\sum_{l=t+1}^{T-1}(\tilde{\vartheta}^{S^{t+1}}_{l})^{\mathbf{t}}\triangle S^{l}\bigg{|}^{2}\bigg{]}$ (3.26)
$\geq\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}t+1,w+\hat{\vartheta}^{\mathbf{t}}_{t}\triangle S^{t},(s^{1:t},S^{t+1})\Big{)}\Big{]}$ (3.27)
$\geq\inf_{\theta\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}t+1,w+\theta^{\mathbf{t}}\triangle S^{t},(s^{1:t},S^{t+1})\Big{)}\Big{]}.$ (3.28)
Here, we use the fact that $\mathcal{F}^{t}_{t}$ is trivial, hence $\hat{\vartheta}_{t}$ is a constant. Using (3.28) repeatedly (for $t\to t,t+1,\cdots,T-1$), we have
$V(t,w,s^{1:t})$ $\geq\inf_{\theta_{t}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}t+1,w+\theta_{t}^{\mathbf{t}}\triangle S^{t},(s^{1:t},S^{t+1})\Big{)}\Big{]}$ (3.29)
$\geq\inf_{\theta_{t}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\inf_{\theta_{t+1}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t},S^{t+1}}}\Big{[}V\Big{(}t+2,w+\theta_{t}^{\mathbf{t}}\triangle S^{t}+\theta_{t+1}^{\mathbf{t}}\triangle S^{t+1},(s^{1:t},S^{t+1},S^{t+2})\Big{)}\Big{]}$ (3.30)
$\geq\inf_{\theta_{t},\theta_{t+1}\in K}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}t+2,w+\theta_{t}^{\mathbf{t}}\triangle S^{t}+\theta_{t+1}^{\mathbf{t}}\triangle S^{t+1},(s^{1:t},S^{t+1},S^{t+2})\Big{)}\Big{]}$ (3.31)
$\geq\cdots$ (3.32)
$\geq\inf_{\theta\in K^{T-t}}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}V\Big{(}T,w+\sum_{l=t}^{T-1}\theta_{l}^{\mathbf{t}}\triangle S^{l},(s^{1:t},S^{t+1:T})\Big{)}\Big{]}$ (3.33)
$=\inf_{\theta\in K^{T-t}}\mathbb{E}_{\mathbb{P}_{s^{1:t}}}\Big{[}\Big{|}1-w-\sum_{l=t}^{T-1}\theta_{l}^{\mathbf{t}}\triangle S^{l}\Big{|}^{2}\Big{]}$ (3.34)
$\geq V(t,w,s^{1:t}).$ (3.35)
The last inequality is due to the fact that constant strategies valued in $K$ are admissible. ∎
3.3 Stability
Denote by $V(\cdot;c,\mathbb{Q})$ the function obtained by DPP (3.23)-(3.24) with $\mathbb{P}$ replaced by $\mathbb{Q}$, and we also write $V=V(\cdot;c,\mathbb{P})$ to emphasize the dependence on $c$ and $\mathbb{P}$. The next theorem aims to bound |VV~||V-$\tilde${V}| by $\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})$.
Theorem 3.8.
For any $\pi\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ and $t=1,2,\cdots,T-1$, we have
$|V(t,w,s^{1:t};c,\mathbb{P})$- $V(t,\tilde{w},\tilde{s}^{1:t};c,\mathbb{Q})|\leq C\Bigg{\{}\big{(}|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|\big{)}\big{(}1+c+|w|+|\tilde{w}|+|s^{t}|+|\tilde{s}^{t}|\big{)}$ (3.36)
$+\big{(}1+c+|w|+|\tilde{w}|+|s^{t}|+|\tilde{s}^{t}|+|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|\big{)}\Big{(}\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+1:T}-\tilde{S}^{t+1:T}|^{2}\Big{)}^{1/2}$ (3.37)
$+\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+1:T}-\tilde{S}^{t+1:T}|^{2}\Bigg{\}}.$ (3.38)
In particular, when $t=1$, we have
$|V(1,0,s^{1};c,\mathbb{P})-V(1,0,s^{1};c,\mathbb{Q})|\leq C\big{\{}(1+s^{1}+c)\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})+\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})^{2}\big{\}}.$ (3.39)
Proof.
For simplicity, we use $V$ and $\tilde{V}$ to denote $V(\cdot;c,\mathbb{P})$ and $V(\cdot;c,\mathbb{Q})$, respectively. We prove (3.38) by induction. $t=T$ is clear from the terminal condition. Suppose we have (3.38) for $t+1$. Because $\pi\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$, we have $\pi_{s^{1:t},\tilde{s}^{1:t}}\in\Pi(\mathbb{P}_{s^{1:t}},\mathbb{Q}_{\tilde{s}^{1:t}})$ (c.f. Backhoff et al. (2017), Proposition 5.1). Therefore, by considering a near-optimal $\vartheta$ in (3.24), it is not hard to show that
$|V(t,w,s^{1:t})-\tilde{V}(t,\tilde{w},\tilde{s}^{1:t})|\leq$ $\sup_{\vartheta\in K}\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\big{|}V(t+1,w+\vartheta^{\mathbf{t}}(S^{t+1}-s^{t}),(s^{1:t},S^{t+1}))$ (3.40)
$-\tilde{V}(t+1,\tilde{w}+\vartheta^{\mathbf{t}}(\tilde{S}^{t+1}-\tilde{s}^{t}),(\tilde{s}^{1:t},\tilde{S}^{t+1}))\big{|}$ (3.41)
Note that in this proof, we adopt the notation $\sim\mathbb{P}, S~\tilde{S}\sim\mathbb{Q}$, and the lower case letters represent realized (sampled) paths, which are used in regular conditional kernels. From Assumption 4, we know $\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+1}|^{2}$ is bounded, uniformly in $s^{1:t}$. Moreover, $\vartheta\in K$ is uniformly bounded by assumption. We thus derive from (3.41) and the induction hypothesis that
$|V(t,w,s^{1:t})-\tilde{V}(t,\tilde{w},\tilde{s}^{1:t})|$ (3.42)
$\leq$ $C\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Bigg{[}\bigg{(}|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|+|S^{t+1}-\tilde{S}^{t+1}|\bigg{)}\bigg{(}1+c+|w|+|\tilde{w}|+|s|+|\tilde{s}|+|S^{t+1}|+|\tilde{S}^{t+1}|\bigg{)}$ (3.43)
$+$ $\bigg{(}1+c+|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|+|S^{t+1}-\tilde{S}^{t+1}|+|w|+|\tilde{w}|+|s|+|\tilde{s}|+|S^{t+1}|+|\tilde{S}^{t+1}|\bigg{)}$ (3.44)
$\times$ $\Big{(}\mathbb{E}_{\pi_{S^{1:t+1},\tilde{S}^{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}+\mathbb{E}_{\pi_{S^{1:t+1},\tilde{S}^{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Bigg{]}$ (3.45)
$=:$ $J_{1}+J_{2}+J_{3}.$ (3.46)
We estimate $J_{i}$ as follows:
$J_{1}$ $\leq C\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Bigg{[}\bigg{(}|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|+|S^{t+1}-\tilde{S}^{t+1}|\bigg{)}\bigg{(}1+c+|w|+|\tilde{w}|+|s^{t}|+|\tilde{s}^{t}|+|S^{t+1}|+|S^{t+1}-\tilde{S}^{t+1}|\bigg{)}\Bigg{]}$ (3.47)
$\leq$ $C\bigg{\{}\big{(}|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|\big{)}\big{(}1+c+|w|+|\tilde{w}|+|s^{t}|+|\tilde{s}^{t}|\big{)}+\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+1}-\tilde{S}^{t+1}|^{2}$ (3.48)
$+\bigg{(}1+c+|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|+|w|+|\tilde{w}|+|s^{t}|+|\tilde{s}^{t}|\bigg{)}\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\big{|}S^{t+1}-\tilde{S}^{t+1}\big{|}$ (3.49)
$+\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\big{|}S^{t+1}-\tilde{S}^{t+1}\big{|}|S^{t+1}|$ (3.50)
$\leq$ $C\bigg{\{}\big{(}|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|\big{)}\big{(}1+c+|w|+|\tilde{w}|+|s^{t}|+|\tilde{s}^{t}|\big{)}+\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+1}-\tilde{S}^{t+1}|^{2}$ (3.51)
$+\bigg{(}1+c+|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|+|w|+|\tilde{w}|+|s^{t}|+|\tilde{s}^{t}|\bigg{)}\bigg{(}\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\big{|}S^{t+1}-\tilde{S}^{t+1}\big{|}^{2}\bigg{)}^{1/2}\bigg{\}}.$ (3.52)
$J_{2}$ $\leq C\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Bigg{[}\bigg{(}1+c+|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|+|w|+|\tilde{w}|+|s|+|\tilde{s}|+|S^{t+1}|+|S^{t+1}-\tilde{S}^{t+1}|\bigg{)}$ (3.53)
$\times\Big{(}\mathbb{E}_{\pi_{S^{1:t+1},\tilde{S}^{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}$ (3.54)
$\leq C\bigg{(}1+c+|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|+|w|+|\tilde{w}|+|s|+|\tilde{s}|\bigg{)}\Big{(}\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}$ (3.55)
$+C\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Bigg{[}\Big{|}S^{t+1}-\tilde{S}^{t+1}\Big{|}\cdot\Big{(}\mathbb{E}_{\pi_{S^{1:t+1},\tilde{S}^{1:t+1}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}\Bigg{]}$ (3.56)
$\leq C\bigg{(}1+c+|w-\tilde{w}|+|s^{t}-\tilde{s}^{t}|+|w|+|\tilde{w}|+|s|+|\tilde{s}|\bigg{)}\Big{(}\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}|S^{t+2:T}-\tilde{S}^{t+2:T}|^{2}\Big{)}^{1/2}$ (3.57)
$+C\Big{(}\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Big{|}S^{t+1}-\tilde{S}^{t+1}\Big{|}^{2}\Big{)}^{1/2}+C\Big{(}\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Big{|}S^{t+2:T}-S^{t+2:T}\Big{|}^{2}\Big{)}^{1/2}.$ (3.58)
$J_{3}=C\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Big{|}S^{t+2:T}-\tilde{S}^{t+2:T}\Big{|}^{2}\leq C\mathbb{E}_{\pi_{s^{1:t},\tilde{s}^{1:t}}}\Big{|}S^{t+1:T}-\tilde{S}^{t+1:T}\Big{|}^{2}$ (3.59)
Combining the bounds of $J_{1}$, $J_{2}$ and $J_{3}$, we establish (3.38). Take $t=1$, $w=\tilde{w}=0$, $\tilde{s}^{1}=s^{1}$ in (3.38), and then take infinimum of all $\pi\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ on the right hand side, we get (3.39) by the definition of $\mathcal{A}\mathcal{W}_{2}$. ∎
Recall the definition of MV value function in (3.2), $v^{*}(\mathbb{P})$. We now provide the stability of $v^{*}$ in terms of the adapted Wasserstein metric.
Corollary 3.9.
Assume $\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})<1$. Then
$|v^{*}(\mathbb{P})-v^{*}(\mathbb{Q})|\leq C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}).$ (3.60)
Proof.
It is not hard to prove a verification result stating that $V(1,0,s^{1};c,\mathbb{Q})=\inf_{\vartheta\in\Theta_{K,\mathbb{Q}}}\mathbb{E}_{\mathbb{Q}}|c-(\vartheta\cdot S)_{T}|^{2}$. Thus, for $\mu=\mathbb{P},\mathbb{Q}$,
$v^{*}(\mu)=\sup_{a\in\mathbb{R}_{+}}\bigg{\{}-\frac{\gamma}{2}V\Big{(}1,0,s^{1};\frac{1}{\gamma}+a,\mu\Big{)}+\frac{1}{2\gamma}+a\bigg{\}}.$ (3.61)
Note that for any $\vartheta\in\Theta_{K,\mathbb{Q}}$,
$\mathbb{E}_{\mathbb{Q}}(\vartheta\cdot S)_{T}$ $\leq\Big{[}\mathbb{E}_{\mathbb{Q}}\sum_{t=0}^{T-1}|\vartheta_{t}||S^{t+1}-S^{t}|\Big{]}$ (3.62)
$\leq L\mathbb{E}_{\mathbb{Q}}|S^{1:T}|$ (3.63)
$\leq L(\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})+\mathbb{E}_{\mathbb{P}}|S^{1:T}|^{2})$ (3.64)
$\leq L.$ (3.65)
Here, $L$ is a generic constant that depends on $T$, $K$ and the bounds of $\mathbb{E}_{\mathbb{P}}[|S^{t+1}|^{2}|\mathcal{F}_{t}]$, and it varies from line to line. Importantly, it does not depends on $\mathbb{Q}$ as long as $\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q})<1$. Consequently, for $\mu=\mathbb{P},\mathbb{Q}$,
$V(1,0,s^{1};c,\mu)\geq c^{2}-2Lc,\quad\forall c>0.$ (3.66)
Therefore, the maximum in (3.61) is attained in a bounded interval $0,\tilde{L}$], and we conclude from (3.39) that
$|v^{*}(\mathbb{P})-v^{*}(\mathbb{Q})|$ $\leq\sup_{a\in[0,\tilde{L}]}\bigg{\{}\frac{\gamma}{2}\Big{|}V\Big{(}1,0,s^{1};\frac{1}{\gamma}+1,\mathbb{P}\Big{)}-V\Big{(}1,0,s^{1};\frac{1}{\gamma}+1,\mathbb{Q}\Big{)}\Big{|}\bigg{\}}$ (3.67)
$\leq C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}),$ (3.68)
completing the proof. ∎
Corollary 3.9 ensures that, measured in adapted Wasserstein metric, MV value functions are stable in the deviation of models. Furthermore, the next corollary justifies that solving the problem with $\mathbb{Q}$ gives a approximated optimal solution to the original problem under $\mathbb{P}$.
Corollary 3.10.
Suppose $\vartheta^{*}_{\varepsilon}$ is a suboptimal solution to $v^{*}(\mathbb{Q})$ such that
$v(\mathbb{Q};\vartheta^{*}_{\varepsilon})\geq v^{*}(\mathbb{Q})-\varepsilon.$ (3.69)
Then, we have
$v(\mathbb{P};\vartheta^{*}_{\varepsilon})\geq v^{*}(\mathbb{P})-\varepsilon-C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}).$ (3.70)
Proof.
In the proof of Theorem 3.8 and Corollary 3.10, we can fixed a $\vartheta^{*}_{\varepsilon}\in\Theta_{K,\mathbb{Q}}$, and because $|\vartheta^{*}_{\varepsilon}|\leq K$, we have
$|v(\mathbb{P};\vartheta^{*}_{\varepsilon})-v(\mathbb{Q};\vartheta^{*}_{\varepsilon})|\leq C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}).$
Therefore, by triangle inequality
$|v(\mathbb{P};\vartheta^{*}_{\varepsilon})-v^{*}(\mathbb{P})|\leq$ $|v(\mathbb{P};\vartheta^{*}_{\varepsilon})-v(\mathbb{Q};\vartheta^{*}_{\varepsilon})|+|v(\mathbb{Q};\vartheta^{*}_{\varepsilon})-v^{*}(\mathbb{Q})|+|v^{*}(\mathbb{Q})-v^{*}(\mathbb{P})|$ (3.71)
$\leq$ $\varepsilon+2C\mathcal{A}\mathcal{W}_{2}(\mathbb{P},\mathbb{Q}),$ (3.72)
which gives (3.70) because $v(\mathbb{P};\vartheta^{*}_{\varepsilon})\leq v^{*}(\mathbb{P})$. ∎
Remark 6.
The method of dynamic programming can also be applied to the fully investment problem, i.e., for admissible set $\Theta_{K,\mathbb{P}}$ we require that $(\theta\cdot S)_{t}=\theta_{t}^{\mathbf{t}}S^{t}$. For the dynamic version, we now take minimum over $\theta\in K(w,s^{t}):=\{\theta^{\mathbf{t}}s^{t}=w\}$, which means that, at each period, we reinvest all of our wealth into risky assets $S$.
4 Applications
In this section, we illustrate the application of diffusion-based adaptive sampling method proposed in Section 2 to MV portfolio selection problems. Specifically, we draw sufficiently many samples from the sampler, and use it as a proxy of the real data model. Results in Section 2 show that these two distributions will be close in $\mathcal{A}\mathcal{W}_{2}$, with sufficiently long diffusion time and sufficiently small score matching error. Meanwhile, from Corollary 3.9 we know, value functions with these two different models will also be close. It thus motivates us to use (deep) reinforcement learning approach, with $\mathbb{Q}$ regarded as an approximated environment, which we call generative environment. Because we can generate large amount of samples with similar statistic properties ($\mathbb{P}\approx\mathbb{Q}$) to the limited amount of data, our approach potentially address the inadequacy and time-sensitivity of financial data.
4.1 Implementation of Algorithm 1
Note that in (2.21) and in Algorithm 1, we need the whole path $X^{1:t}$ (or $s^{1:t}$ in the context of portfolio selection problems) as inputs in both training and sampling procedure to account for non-Markovian dependency in time-series data. This makes our algorithm suffer from curse of dimensionality. To resolve this, we use the feature extraction trick. More specifically, inspired by some papers on conditional diffusion model, such as Rasul et al. (2021) and Yan et al. (2021), we suppose there exists a feature extraction function with varying length of input (which is realized as an RNN encoder), $R:\mathbb{R}^{td}\to\mathbb{R}^{d^{\prime}}$, where $t=1,2,\cdots,T$ and $d^{\prime}$ is the dimension of feature space. Given any path $s^{1:T}$, we define hidden variables $h^{1:T}$ by $h^{t}=R(s^{t},h^{t-1})$, and suppose that $s$ can be recovered from $h$: $s^{1:t}=(s^{1:t})^{\prime}$ if and only if $h^{t}=(h^{t})^{\prime}$. Under this assumption, we propose to train the RNN encoder $R_{\theta}:\mathbb{R}^{d}\times\mathbb{R}^{d^{\prime}}\to\mathbb{R}^{d^{\prime}}$ and score network $s_{\theta}:[0,\mathcal{T}]\times\mathbb{R}^{d^{\prime}}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ jointly using the following variant of (2.21):
$\frac{1}{TM}\sum_{t,m}\Big{|}\sqrt{1-e^{2\tau_{(m)}}}s_{\theta}\Big{(}\tau_{(m)},h^{t}_{(m)},s^{t}_{(m)}e^{-\tau_{(m)}}+\sqrt{1-e^{-2\tau_{(m)}}}\mathbf{z}_{(m)}\Big{)}+\mathbf{z}_{(m)}\Big{|}^{2}.$ (4.1)
Here, $\{s^{1:T}_{(m)}\}_{m=1}^{M}$ is a batch of data, and $\{h^{1:T}_{(m)}\}_{m=1}^{M}$ are the corresponding hidden variables obtained by feeding the data to the RNN encoder $R_{\theta}$; $\{\tau_{(m)}\}_{m=1}^{M}$ are randomly sampled from $0,\mathcal{T}]$; $\{\mathbf{z}_{(m)}\}_{m=1}^{M}$ are sampled from pure noise $\mathcal{N}(0,I)$. Although only $s_{\theta}$ explicitly appears in (4.1), $R_{\theta}$ is implicitly trained by backward propagation. Algorithm 2 provides the pseudocode for the training procedure.
Algorithm 2 Training the score network and the RNN encoder
Initialization: Initialize $s_{\theta}$ and $R_{\theta}$; collect the training data.
1:for epochs do
2:  Sample a batch of data $\{s^{1:T}_{(m)}\}_{m=1}^{M}$.
3:  Compute $\{h^{1:T}_{(m)}\}_{m=1}^{M}$ using $R_{\theta}$.
[…]
5:  Sample $\{\mathbf{z}_{(m)}\}_{m=1}^{M}$ from $\mathcal{N}(0,I)$.
6:  Update $\theta$ by minimizing (4.1).
7:end for
Outputs: Trained score network $s_{\theta}$ and RNN encoder $R_{\theta}$.
As for the sampling procedure, we consider an implementable version of Algorithm 1, which has the following two refinements: we utilize the hidden variables obtained from the RNN encoder to resolve high dimensionality; we also use the conventional predictor-corrector sampler to reduce the discretization error from the SDE solver. The sampling procedure can be summarized in Algorithm 3.
Algorithm 3 Adaptive sampling: implementation
Inputs: Trained score network $s_{\theta}$ and RNN encoder $R_{\theta}$; number time discretization (predictor) steps $\mathrm{N_{pre}}$; number of corrector steps $\mathrm{N_{cor}}$; predictor steps $\{\tau_{k}\}_{k=1}^{\mathrm{N_{pre}}}$; corrector step sizes $\{\mathbf{e}_{k}\}_{k=1}^{\mathrm{N_{pre}}}$; initial feature $h^{0}$
[…]
2: Sample $s^{1:T}\sim\mathcal{N}(0_{\mathbb{R}^{dT}},I_{dT\times dT})$
3:for $t=1,\cdots,T$ do
4:  for $k=0,1,\cdots,\mathrm{N_{pre}}-1$ do
5:    $s^{t}\leftarrow s^{t}+\Big{(}s^{t}+2s_{\theta}\big{(}\mathcal{T}-\tau_{k},h,s^{t}\big{)}\Big{)}\cdot(\tau_{k+1}-\tau_{k})+\sqrt{2}\bm{\varepsilon}$, where $\bm{\varepsilon}\sim\mathcal{N}(0,t_{k+1}-t_{k})$.
6:   # The predictor step.
7:   for $l=0,1,\cdots,\mathrm{N_{cor}}-1$ do
8:     $s^{t}\leftarrow s^{t}+\mathbf{e}_{k}s_{\theta}\big{(}\tau_{k},h,s^{t}\big{)}+\sqrt{2}\bm{\varepsilon}^{\prime}$, where $\bm{\varepsilon}^{\prime}\sim\mathcal{N}(0,\mathbf{e}_{k})$.
9:    # The corrector step.
10:   end for
11:  end for
12:   $h\leftarrow R_{\theta}(s^{t},h)$.  #Compute $h^{t}$ from $h^{t-1}$.
13:end for
Outputs: $s^{1:T}$.
4.2 Solving portfolio choice problem with $\mathbb{Q}$
In this subsection, we employ our pretrained generative model, $\mathbb{Q}$, as a scenario generator. Specifically, we propose to utilize reinforcement learning within a generative environment provided by $\mathbb{Q}$. See Algorithm 4. Note that this is a refined version of the deterministic policy gradient algorithm (DDPG, c.f. Lillicrap et al. (2016)). We redesign part of the algorithm to accommodate the unique characteristics of portfolio selection problems. These adjustments include separating the scenario pool and replay buffer, which facilitates the convenient integration of the generative model.
For reinforcement learning in financial applications, it is advisable to refrain from utilizing long-term data, as certain global factors, such as macroeconomic conditions or policy changes, significantly alter the patterns of financial data, rendering the learned model unsuitable for practical implementation. However, a limited time period results in a scarcity of data, diminishing the performance of the algorithm. With our approach, this issue is addressed by generating a sufficient quantity of data from $\mathbb{Q}$, which closely approximates the current real model, $\mathbb{P}$. Additionally, we can retrain $\mathbb{Q}$ with certain schedule to enable long-time implementation of our algorithm (see an example of retraining trick in Subsection 4.3.2).
Algorithm 4 Policy gradient with generative environment
Initialization: Initialize Q-network $Q_{\alpha}$ and policy network $\pi_{\beta}$; initialize target networks: $\alpha_{\mathrm{targ}}\leftarrow\alpha$, $\beta_{\mathrm{targ}}\leftarrow\beta$; set up an empty replay buffer $\mathcal{D}$ and an empty scenario pool $\mathcal{S}$, both with size $L$; train the score network $s_{\theta}$ and RNN encoder $R_{\theta}$ using Algorithm 2. Collect context window data $s^{-T:-1}$.
1:  Use $s^{-T:-1}$ to compute initial feature $h^{0}$.
2:  Sample $L$ paths prediction $s^{1:T}$ and compute corresponding feature paths $h^{1:T}$ using Algorithm 3; store them in $\mathcal{S}$
3:  for steps  do
4:   Initialize $w=1$; sample one path $(s^{1:T},h^{1:T})$ from $\mathcal{S}$; sample $c$ from Exp(10).
5:   for $t=1$ to T1T-1 do
6:    if warm-up  then
7:     $a\sim$ Uniform(K)
8:    else
9:     $a\leftarrow\mathcal{P}_{K}(\pi_{\beta}(t,w,h^{t},c)+\varepsilon)$.
10:    end if
11:    Store $(w,a,c)$ in $\mathcal{D}$.
12:    $w\leftarrow w+a\cdot(s^{t+1}-s^{t})$.
13:    if not warm-up then
14:      Sample a batch $\{(s^{1:t+1}_{(j)},h^{1:t+1}_{(j)}\}_{j=1}^{B}$ from $\mathcal{S}$ and $\{(w_{(j)},a_{(j)},c_{(j)}\}_{j=1}^{B}$ from $\mathcal{D}$.
15:      $w^{\prime}_{(j)}=w_{(j)}+a_{(j)}\cdot(s_{(j)}^{t+1}-s_{(j)}^{t}).$
16:     if $t\leq T-1$ then
17:       Update $\alpha$ by minimizing
$\frac{1}{B}\sum_{j}\Big{|}Q_{\alpha}(t,w_{(j)},h_{(j)}^{t},a_{(j)},c_{(j)})-Q_{\alpha_{\mathrm{targ}}}\Big{(}t+1,w^{\prime}_{(j)},h_{(j)}^{t+1},\pi_{\beta_{\mathrm{targ}}}(t+1,w^{\prime}_{(j)},h^{t+1}_{j},c_{(j)}),c_{(j)}\Big{)}\Big{|}^{2}.$
18:     else
19:      Update $\alpha$ by minimizing
$\frac{1}{B}\sum_{j}\Big{|}Q_{\alpha}(t,w_{(j)},h^{t}_{(j)},a_{(j)},c_{(j)})-|w_{(j)}-c_{(j)}|^{2}\Big{|}^{2}.$
20:     end if
21:      Update $\beta$ by gradient ascent via
$\frac{1}{B}\sum_{j}\nabla_{a}Q_{\alpha}\Big{(}t,w_{(j)},h_{(j)}^{t},\pi_{\beta}(t,w_{(j)},h_{(j)}^{t},c_{(j)}),c_{(j)}\Big{)}\nabla_{\beta}\pi_{\beta}(t,w_{(j)},h^{t}_{(j)},c_{(j)})$
22:      Update target network:
$\alpha_{\mathrm{targ}}\leftarrow\rho\alpha_{\mathrm{targ}}+(1-\rho)\alpha,$ (4.2)
$\beta_{\mathrm{targ}}\leftarrow\rho\beta_{\mathrm{targ}}+(1-\rho)\beta.$ (4.3)
23:    end if
24:   end for
25:  end for
26:  Find the optimal multiplier by maximizing
$\frac{\gamma}{2}Q_{\alpha}\big{(}1,0,h^{1},\pi_{\beta}(1,0,h^{1},c),c\big{)}+c.$
Output: Policy network $\pi_{\beta}$
4.3 Numerical experiments
This section provides numerical experiment results to showcase the feasibility of proposed algorithms, including the adpative sampling through diffusion model, and the policy gradient algorithm to solve portfolio selection problems. We conduct both sythetic data and real data experiments. All codes are available in https://github.com/fyyuan/diffusion_dynamic_mv.
4.3.1 Synthetic data
To test the feasibility of reinforcement learning part of our proposed algorithm, in this section we implement Algorithm 4 in a simulated environment. Specifically, we generate return and price data from a known distribution, collect these data, and utilize them for both training and testing purposes. The details of the simulation are as follows: we set $T=12$ and $d=10$ (i.e., we consider 10 assets and make portfolio decisions at $t=1,2,\cdots,12$). We suppose the prices satisfies a simple normal return model, i.e., $\{r_{t}\}_{t=1}^{12}$ are i.i.d. from $\mathcal{N}(\mu/12,\Sigma/12)$, with
$\mu=(0.15,0.12,0.18,0.14,0.25,0.10,0.08,0.09,0.07,0.11)^{\mathbf{t}},$ (4.4)
$\Sigma=\left(\begin{array}[]{cccccccccc}0.1024&0.0627&0.1008&0.0672&0.1568&0.0320&0.0230&0.0256&0.0448&0.0282\\ 0.0627&0.0784&0.0882&0.0588&0.1372&0.0280&0.0202&0.0224&0.0392&0.0246\\ 0.1008&0.0882&0.2025&0.0945&0.2205&0.0450&0.0324&0.0360&0.0630&0.0396\\ 0.0672&0.0588&0.0945&0.0900&0.1470&0.0300&0.0216&0.0240&0.0420&0.0264\\ 0.1568&0.1372&0.2205&0.1470&0.4900&0.0700&0.0504&0.0560&0.0980&0.0616\\ 0.0320&0.0280&0.0450&0.0300&0.0700&0.0625&0.0135&0.0150&0.0350&0.0330\\ 0.0230&0.0202&0.0324&0.0216&0.0504&0.0135&0.0324&0.0288&0.0252&0.0119\\ 0.0256&0.0224&0.0360&0.0240&0.0560&0.0150&0.0288&0.0400&0.0280&0.0132\\ 0.0448&0.0392&0.0630&0.0420&0.0980&0.0350&0.0252&0.0280&0.1225&0.0231\\ 0.0282&0.0246&0.0396&0.0264&0.0616&0.0330&0.0119&0.0132&0.0231&0.0484\end{array}\right)$ (4.5)
Here, to fit the simulated data real-world scale, we randomly generate the annualized return and covariance according to the typical range, and rescale them so that $T$ represents one year. In the simulation study, the scenario pool $\mathcal{S}$ is constructed by setting $s^{0}\sim{\rm Uniform}(1,200)$ and generated the path $s^{1:T}$ by sampling returns independently from $\mathcal{N}(\mu/12,\Sigma/12)$. We then train the actor and critic networks with Algorithm 4. To enhance the performance and stability, we use the twin delayed version of Algorithm 4 (Twin Delayed DDPG, TD3; c.f. Fujimoto et al. (2018)); we refer interested readers to our github repository for details. Here, we simply use the price as features, i.e., $h\equiv s$. We consider a constrained portfolio selection problem, in which both short-selling and borrowing are prohibited. In this case, we set the output of actor network to be the proportion of total wealth invested in each stock. Therefore, the action space is $\triangle_{d-1}:=\{a=(a^{1},\cdots,a^{d}):a^{i}\geq 0,\sum_{i}a^{i}=1\}$.
In the testing phase, we set an extended time horizon $T_{\rm test}=10T$ and deploy the trained agent to make portfolio decision for $t=1,2,\cdots,T_{\rm test}$. We generate $L_{\rm test}=100$ testing scenarios, independently from training scenario pool. For each testing scenario and any portfolio strategy, we can divide the whole testing horizon into 10 testing periods (years) and record returns in each period. Then, the (mean) return, volatility and Sharpe ratio are calculated for each testing scenario. We finally report mean and standard variation of these three metrics across all testing periods. Choosing $\gamma=1.5$, for each testing period we can also compute the mean variance data using results of different testing scenario, and report the averaged mean variance for 10 periods. To ensure results of different testing periods are in the same scale, we compute the mean variance of return rate instead of the net gain.
We compare three portfolio strategies:
  1. 1.
    Equal Weight: $a\equiv(1/d,1/d,\cdots,1/d)$;
  2. 2.
    Static Markowitz: We first solve Markowitz problem (with no shortselling, no borrowing and full investment constraints) with $\mu$ and $\Sigma$ and denote by $a_{\rm mk}$ the solution. Then, at the beginning of each testing period, we adjust the position so that total wealth is allocated according to $a_{\rm mk}$ and hold for the testing period.
  3. 3.
    TD3: For each $t=1,2,\cdots,T_{\rm test}$, then action $a_{t}$ is given by the actor network trained by Algorithm 4. When $t>T$, for numerical stability, we normalize the feature scale to the first test period by $h^{t}=s^{0}s^{t}/s^{t_{0}}$ if $t=kT+t_{0}$, $k=0,1,\cdots,9$.
Training and testing are conducted on a personal Macbook Air. The training is efficient with only 379s running time. In Figure 3, we present the training log, i.e., the performance of the agent after each training epoch, where one epoch has 50 steps. For comparison, we use the same testing scenario for every epoch. In Table 1, we report the evaluation results after a total of 10,000 steps of training. To better illustrate the performance of the agent after training, we also present the portfolio decision and wealth trajectory in one testing scenario in Figure 4.
(a) Return rate
(b) Sharpe ratio
(c) Mean variance value
Figure 3: Testing results after each epoch of training.
Table 1: Testing result of the final output
Return Volatility Sharpe MV
Equal Weight 12.96%(0.08) 0.25(0.07) 0.51(0.31) 0.07
Static Markowitz 13.92%(0.08) 0.24(0.07) 0.57(0.32) 0.09
TD3 13.01%(0.08) 0.23(0.05) 0.61(0.36) 0.09
Figure 4: Agent performance in one testing scenario
From Figure 3 and Table 1, we argue that Algorithm 4 obtains an effective portfolio strategy. Following low-cost training, it performs better than the equal weight strategy, achieving higher returns, Sharpe ratios, and mean-variance values. It is widely acknowledged in literature that the equal weight strategy serves as a crucial baseline for portfolio evaluation. Surprisingly, outputs of many model-based algorithms do not consistently performance better than the equal weight strategy. For instance, see DeMiguel et al. (2009), Blanchet et al. (2022), Huang et al. (2024b). Furthermore, the agent also acheives comparable returns and mean variance values, while also exhibiting better volatility and Sharpe ratios compared to the static Markowitz strategy. The Markowitz strategy possesses complete knowledge of the model in the simulation study and fully utilizes the model information. Consequently, achieving comparable or superior results than Markowitz implies that the agent uncovers hidden information (such as return and volatility) solely from price data.
4.3.2 Real data
In the previous subsection, we showcase the performance of the reinforcement learning part of Algorithm 4. In other words, the scenario pool $\mathcal{S}$ is sampled from a known distribution. The ability to sample many paths from this distribution is crucial to the successful implementation of Algorithm 4. In this section, we install the diffusion model in Algorithm 4 as a sampler of the path distribution that has similar distributional properties to the real data. Although the algorithm has theoretical guarantee by Theorems 2.3 and 3.8, we still need to justify its applicability to real-world scenario. Our empirical tests employ the value-weighted 10 Industry Portfolios from the Kenneth R. French Data Library, a monthly panel that aggregates all common stocks on the CRSP tape into ten broad industry groups—Non-Durables (’NoDur’), Durables (’Durbl’), Manufacturing (’Manuf’), Energy (’Enrgy’), HiTech (’HiTec’), Telecommunications (’Telcm’), Utilities (’Utils’), Shops (’Shops’), Health (’Hlth’), and Others (’Other’). We use the series of excess returns (portfolio return minus the one-month Treasury bill rate) spanning July 1926 through March 2025. This dateset is splitted as follows:
  1. 1.
    The test dataset comprises data collected over the past 16 years, from April 2009 to March 2025. The initial year of the test dataset, spanning from April 2009 to March 2008, serves as the historical window and is utilized as the initial condition input into the diffusion model; see variable $h$ in Algorithms 2-4. Consequently, the evaluation period is from March 2010 to March 2025.
  2. 2.
    The validation dataset consists of data from August 1992 to March 2009, encompassing 500 months prior to the test dataset.
  3. 3.
    The remaining dataset is employed in Algorithm 2 to train the diffusion model $s_{\theta}$ and $R_{\theta}$.
For training, we use a same‐resolution UNet which is inspired by Ronneberger et al. (2015) as the score network $s_{\theta}$ and an LSTM Hochreiter and Schmidhuber (1997) as the RNN encoder $R_{\theta}$. To improve the performnance, we employ the refined version of OU process: the variance‐preserving stochastic differential equation (VPSDE):
$\mathrm{d}X_{\tau}=-\tfrac{1}{2}\,\beta(\tau)\,X_{\tau}\,\mathrm{d}\tau+\sqrt{\beta(t)}\,\mathrm{d}B_{\tau},\qquad\beta(\tau)=\beta_{\min}+(\beta_{\max}-\beta_{\min})\,\tau,$
discretised into $N$ uniform timesteps $\{t_{n}\}_{n=1}^{N}$. Default hyperparameters are $\beta_{\min}=0.01$, $\beta_{\max}=10.0$, and $N=1000$. Then, $s_{\theta}$ and $R_{\theta}$ are trained according to Algorithm 2. Detailed model architectures and training configurations are avaible in our github repository. Optimisation uses the AdamW optimiser with learning rate $10^{-3}$, $(\beta_{1},\beta_{2})=(0.9,0.999)$, and weight decay $10^{-2}$. A cosine‐annealing scheduler runs for 400 epochs with a 5 epoch warm-up. Gradients are clipped to a maximum norm of 1, and mixed precision training (bfloat16) is enabled via torch.cuda.amp. We maintain an exponential moving average (EMA) of model weights with decay 0.999 and evaluate validation metrics on the EMA weights. With these settings, training on a single NVIDIA T-40 is completed in about 10 minutes.
For testing, we set the time horizon $T=12$ (i.e., one year) for TD3 agent, and use an extended test time horizon $T_{\rm test}=15T$ (i.e., 15 years) to show the robustness. We compare the market benchmark, the S&P 500 index, and four portfolio strategies with details described below:
  1. 1.
    Equal Weight (EW): $a\equiv(1/d,1/d,\cdots,1/d)$;
  2. 2.
    History-based Markotwitz (HistMarkowitz): For each $t=1,2,\cdots,T_{\rm test}$, action $a$ is determined by solving the Markowitz problem with no short selling, no borrowing and fully investment constraints. The mean and covariance are estimated by the sample mean and sample covariance of the most recent 60 months data (as in DeMiguel et al. (2009)). In other words, this is a monthly re-balancing Markowitz strategy in which we use historical data as estimators;
  3. 3.
    Generative-model-based Markowitz (GenMarkowitz): For each $t=1,2,\cdots,T_{\rm test}$, we collect the most recent $T$ months (i.e., one year) data as the context window, and use it as the inital feature $h^{0}$ in the diffusion model (see Algorithm 3). Then, we sample 500 samples of one month predictions for $s^{t}$. Then, we use this prediction to solve constrained Markowitz problem. In other words, this is a monthly re-balancing Markowitz strategy based on generative model;
  4. 4.
    Generative-model-based TD3 (GenTD3): We set the first year returns from the test dataset as the context window $s^{-T:-1}$}, and then adaptively sample 500 paths of one-year prediction $s^{1:T}$ based on Algorithm 3. We use these samples as the scenario pool $\mathcal{S}$ to train the TD3 agent. The trained agent is used repeatedly for the first 7.5 years, outputing the portfolio with updated feature $h^{t}$ for any $t=1,2,\cdots,T_{\rm Test}$. Then, at the midway of the testing period, we use the most recent one-year data as a new context window and generate another set of 500 paths. The agent is retrained using these paths and then used for the rest 7.5 years [^2] We only retrain the model once due to the limited computation resources. Ideally, retraining the RL agent each year, which is aligned with the length of test period, leads the agent better utilized the up-to-date market model. However, this simple retraining scheme presented here has been sufficient to illustrate the application of our theories..
Remark 7.
For clarity, we emphasize that in the GenTD3 strategy, we evaluate three neural networks: the score network $s_{\theta}$, the RNN encoder $R_{\theta}$ and the policy network $\pi_{\beta}$. They are all coupled through the training phase. See Figure 2 in the Introduction part.
We set $\gamma=3$ and test the four strategies comprehensively using the following metrics: annualized return, annualized volatility, Sharpe ratio, Sortino ratio, maximal drawdown and Calmer ratio. See Table 2 for the performance. We also report the output portfolio for GenTD3 strategy and the wealth trajectory for all four strategies; see Figures 5. From the output tables and figures, we conclude that two generative-model-based strategies outperform S&P 500 and two baselines (in terms of Sharpe and Sortino ratio), justifying the efficiency of the diffusion model. We emphasize that, as a time-series predictor, the diffusion model consists of two parts: the score network $s_{\theta}$ and the RNN encoder $R_{\theta}$. They are both trained in an adaptive manner via the proposed Algorithm 2. For the GenMarkowitz strategy, while the feature variables $h^{t}$ are updated online, it treats the portfolio selection as a one-period problem and solves static Markowitz problems for every $t$. This is in contrast to GenTD3 strategy, which deal with a dynamic decision problem. We observe from Table 2 that the GenTD3 strategy outperforms GenMarkowitz in 5 out of 6 metrics, except annualized return. These results support our proposed adaptive sampling method (Algorithm 3), which includes intertemporal structure in predictions. The GenTD3 agent then exploits such intertemporal structure through an reinforcement learning framework (Algorithm 4).
Table 2: Performance of different portfolio strategies with $\gamma=3$
Return Volatility Sharpe Sortino Max Drawdown Calmer
S&P 500 10.98% 0.1460 0.7522 0.695519 -0.2477 0.4433
EW 13.26% 0.1502 0.8831 0.8310 -0.2293 0.5783
HistMarkowitz 8.08% 0.1553 0.5200 0.5226 -0.3014 0.2679
GenMarkowitz 13.59% 0.1536 0.8850 0.8526 -0.2690 0.5053
GenTD3 12.80% 0.1398 0.9158 0.8966 -0.2325 0.5507
Figure 5: Output portfolio and wealth trajectory
Because it is difficult to specify a level of risk aversion in reality, we need to justify the robustness of our algorithm under the change of $\gamma$. We report the results in Tables 3-4 for a lower risk aversion $\gamma=0.5$ and a higher risk aversion $\gamma=5$. We observe that, although GenMarkwotiz performs exceptional for $\gamma=5$, it crashes for smaller $\gamma=0.5$, even underperforming EW and S&P 500. GenTD3, in contrast, is stable in varying risk aversion and achieves the best results for $\gamma=0.5$. It also has a performance comparable to GenMarkowitz for $\gamma=5$, better than the benchmarks. This sensitivity analysis indicates the importance of the proposed adaptive sampling scheme. We also conclude that, the two strategies based on diffusion model are suitable for different risk aversions. In particular, choosing the appropriate risk aversion (which is not an easy task, though), GenMarkowitz may also have satisfactory performance, justifying the applicability of diffusion models in financial data.
Table 3: Performance of different portfolio strategies with $\gamma=0.5$
Return Volatility Sharpe Sortino Max Drawdown Calmer
S&P 500 10.98% 0.1460 0.7522 0.6955 -0.2477 0.4433
EW 13.26% 0.1502 0.8831 0.8310 -0.2293 0.5783
HistMarkowitz 6.18% 0.2043 0.3025 0.2973 -0.4090 0.1511
GenMarkowitz 12.02% 0.2418 0.4970 0.5121 -0.5563 0.2160
GenTD3 13.57% 0.1440 0.9428 0.9774 -0.2199 0.6405
Table 4: Performance of different portfolio strategies with $\gamma=5$
Return Volatility Sharpe Sortino Max Drawdown Calmer
S&P 500 10.98% 0.1460 0.7522 0.6955 -0.2477 0.4433
EW 13.26% 0.1502 0.8831 0.8310 -0.2293 0.5783
HistMarkowitz 9.26% 0.1375 0.6737 0.6909 -0.2605 0.3555
GenMarkowitz 14.17% 0.1360 1.0418 1.0543 -0.1838 0.7711
GenTD3 12.73% 0.1392 0.9150 0.8916 -0.2326 0.5474
5 Concluding remarks
In this work, we extend the score-based diffusion‑model framework to handle time‑series data. We introduce an adaptive training and sampling scheme tailored to dynamic control problems, and we demonstrate its effectiveness on the dynamic mean–variance portfolio‑selection problem. Our theoretical contributions include a quantification‑error bound and a model‑stability result that make this application feasible. We implement the proposed algorithms and present experiments that showcase the efficiency of the approach.
Several future directions naturally emerge. In the theoretical aspect, one could relax the assumptions on the data distribution—particularly Assumption 2. With a more refined approximation analysis, it may be possible to remove Lipschitz continuity in the conditional variable $x^{1:t}$ and impose it only on $x^{t+1}$ in (2.25), which aligns with the usual Lipschitz condition for the score function. Moreover, the error bounds established here are unlikely to be optimal; tightening, or even optimizing, the rate remains an interesting direction. Practically, further fine‑tuning of the proposed algorithms and large‑scale experiments are highly motivated. Furthermore, to enhance the predictive ability we can incorporate more covariant variables into the input of RNN encoder, instead of just using the price (return) data. Quantitative analysis of the RL part of this work (e,g, the regret analysis) is also interesting. We leave these directions for future study.
Appendix A On Assumption 3
In this appendix, we demonstrate that the approximating network $s_{\theta}$ can be selected to fulfill Assumption 3 while simultaneously possessing the desired approximation power, i.e., satisfying Assumption 1. For technical reasons, presented in this appendix, we provide a proof that Assumption 1 can be achieved with $\tau\in[\tau_{0},\mathcal{T}]$, where $\tau_{0}>0$. By using the standard early stopping technique employed in the diffusion model literature, we still obtain a convergence guarantee by error bounds in Theorem 2.3. To avoid detracting from the central focus of our discussion, in the main context we directly assume that Assumption 1 can be made uniform for $\tau\in[0,\mathcal{T}]$.
To provide desired score matching error (2.3) and (2.4) in a unified manner, we work on the pair $(x^{1:t-1},x^{t})$ for $t\in\{1,2,\cdots,T\}$, where $x^{1:t-1}$ is the condition (observation) and $x^{t}$ is the prediction of the next time stamp. If $t=1$, then $x^{1:0}$ can be regarded as a void variable and $\mathbb{P}_{x^{1:0}}=\mathbb{P}_{1}$. This corresponds to the initial score matching step (i.e., (2.3)).
To construct our approximating network, we first observe that, by (2.7), the real score function $\nabla_{x}\log p_{t}(\tau,x|x^{1:t-1})$ can be written as
$\nabla_{x}\log p_{t}(\tau,x|x^{1:t-1})$ $=\frac{\nabla_{x}\int_{\mathbb{R}^{d}}\phi(\tau,x|x^{t})\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}{\int_{\mathbb{R}^{d}}\phi(\tau,x|x^{t})\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}$ (A.1)
$=-\frac{\int_{\mathbb{R}^{d}}(x-x^{t}h_{1}(\tau))e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}{h_{2}(\tau)\int_{\mathbb{R}^{d}}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}$ (A.2)
$=-\frac{x}{h_{2}(\tau)}+\frac{h_{1}(\tau)}{h_{2}(\tau)}\frac{\int_{\mathbb{R}^{d}}x^{t}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}{\int_{\mathbb{R}^{d}}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}$ (A.3)
$=:-\frac{x}{h_{2}(\tau)}+G(\tau,x^{1:t-1},x).$ (A.4)
Here, for simplicity we denote $h_{1}(\tau):=e^{-\tau}$, $h_{2}(\tau):=1-e^{-2\tau}$. This decomposition motivates us to find an appropriate $G_{\theta}$ to approximate $G$, and define
$s_{\theta}(\tau,x^{1:t-1},x):=-\frac{x}{h_{2}(\tau)}+G_{\theta}(\tau,x^{1:t-1},x).$ (A.5)
If $G_{\theta}$ can be chosen such that it is Lipschitz continuous and compactly supported in $\{|x^{1:t}|<R\}$, then Assumption 3 is satisfied with $M_{\rm disp}\sim R^{2}$. Indeed, noting that $h_{2}(\tau)<1$, we have
$2x\cdot s_{\theta}(\tau,x^{1:t-1},x)$ $=-\frac{2|x|^{2}}{h_{2}(\tau)}+2x\cdot G_{\theta}(\tau,x^{1:t-1},x)$ (A.6)
$\leq-2|x|^{2}+M|x|$ (A.7)
$\leq-(1+\delta)|x|^{2}+\frac{M^{2}}{4(1-\delta)},$ (A.8)
where $M$ is the upper bound of $G_{\theta}$. Because the real score function, hence $G$, is assumed to be Lipchitz (say, with a Lipchitz constant $L$), we have $M\leq\|G-G_{\theta}\|_{\infty}+\sup_{|x^{1:t}|<R}|G(\tau,x^{1:t})|\leq 1+LR$. This is exactly Assumption 3 with $M_{\rm disp}\sim R^{2}$.
In fact, we can indeed construct a bounded supported $G_{\theta}$ which approximate $G$ with any accuracy $\varepsilon_{\rm score}$, with the support size $R$ satisfies $\varepsilon_{\rm score}\sim e^{-cR^{2}}$. To prove this, we need an improved version of universal approximation property of the ReLU network from Chen et al. (2022):
Lemma A.1 (Lemma 10 in Chen et al. (2022)).
For any $\varepsilon>0$, there exists a ReLU network architecture $\hat{f}$ on $\mathbb{R}^{m}$ such that for any 1-Lipschitz function $f$ defined on $0,1]^{m}$ with $f(0)=0$, we have $\|f-\hat{f}\|_{\infty}\leq\varepsilon$. Moreover, $\hat{f}$ is $10d$-Lipschitz.
Lemma A.2.
For any $\varepsilon_{\rm score}>0$, there exists a constant $R>0$ and a ReLU network $G_{\theta}$, such that $G_{\theta}(\tau,\cdot,\cdot)$ is supported on $B_{R}:=\{|x^{1:t-1}|\leq R\}\times\{|x|\leq R\}$ for any $\tau\in[\tau_{0},\mathcal{T}]$, where $e^{-cR^{2}}\sim\varepsilon_{\rm score}$ and
$\sup_{\tau\in[\tau_{0},\mathcal{T}]}\mathbb{E}_{X^{1:t-1}\sim\mathbb{P}_{1:t-1}}\mathbb{E}_{X_{\tau}^{t}\sim p_{t}(\tau,\cdot|X^{1:t-1})}\bigg{|}G_{\theta}(\tau,X^{1:t-1},X^{t}_{\tau})-G(\tau,X^{1:t-1},X^{t}_{\tau})\bigg{|}^{2}\leq\varepsilon^{2}_{\rm score}.$ (A.9)
Proof.
We first observe that
$\mathbb{E}_{X^{1:t-1}\sim\mathbb{P}_{1:t-1}}\mathbb{E}_{X_{\tau}^{t}\sim p_{t}(\tau,\cdot|X^{1:t-1})}\Bigg{[}\bigg{|}G_{\theta}(\tau,X^{1:t-1},X^{t}_{\tau})-G(\tau,X^{1:t-1},X^{t}_{\tau})\bigg{|}^{2}I_{B_{R}^{c}}\Bigg{]}$ (A.10)
$=$ $\mathbb{E}_{X^{1:t-1}\sim\mathbb{P}_{1:t-1}}\mathbb{E}_{X_{\tau}^{t}\sim p_{t}(\tau,\cdot|X^{1:t-1})}\Bigg{[}\bigg{|}G(\tau,X^{1:t-1},X^{t}_{\tau})\bigg{|}^{2}I_{B_{R}^{c}}\Bigg{]}$ (A.11)
$=$ $\frac{h_{1}(\tau)}{h_{2}(\tau)}\int_{(x^{1:t-1},x)\in B_{R}^{c}}\Bigg{|}\frac{\int_{\mathbb{R}^{d}}x^{t}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}{\int_{\mathbb{R}^{d}}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}\Bigg{|}^{2}p_{t}(\tau,x|x^{1:t-1})\mathrm{d}x\mathbb{P}_{1:t-1}(\mathrm{d}x^{1:t-1})$ (A.12)
$\leq$ $\frac{h_{1}(\tau)}{h_{2}(\tau)}\int_{(x^{1:t-1},x)\in B_{R}^{c}}\frac{\int_{\mathbb{R}^{d}}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}{\int_{\mathbb{R}^{d}}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})}p_{t}(\tau,x|x^{1:t-1})\mathrm{d}x\mathbb{P}_{1:t-1}(\mathrm{d}x^{1:t-1})$ (A.13)
$=$ $(2\pi h_{2}(\tau))^{-d/2}\frac{h_{1}(\tau)}{h_{2}(\tau)}\int_{(x^{t-1},x)\in B_{R}^{c}}\int_{\mathbb{R}^{d}}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x\mathbb{P}_{1:t-1}(\mathrm{d}x^{1:t-1}).$ (A.14)
Here, we use (2.7). Clearly, $B_{R}^{c}\subset\{|x^{1:t-1}|>R\}\cup\{|x|>R\}$. Thus by splitting the integrals in (A.14) we have
$\mathbb{E}_{X^{1:t-1}\sim\mathbb{P}_{1:t-1}}\mathbb{E}_{X_{\tau}^{t}\sim p_{t}(\tau,\cdot|X^{1:t-1})}\Bigg{[}\bigg{|}G_{\theta}(\tau,X^{1:t-1},X^{t}_{\tau})-G(\tau,X^{1:t-1},X^{t}_{\tau})\bigg{|}^{2}I_{B_{R}^{c}}\Bigg{]}$ (A.15)
$\leq$ $C\Bigg{(}\int_{|x^{1:t-1}|\geq R}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x\mathbb{P}_{1:t-1}(\mathrm{d}x^{1:t-1})$ (A.16)
$+\int_{\mathbb{R}^{d(t-1)}}\int_{|x|>R}\int_{\mathbb{R}^{d}}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x\mathbb{P}_{1:t-1}(\mathrm{d}x^{1:t-1})\Bigg{)}$ (A.17)
$=:I_{1}(R)+I_{2}(R),$ (A.18)
where $C$ is a constant not depending on $R$, $\theta$ and $\varepsilon_{\rm score}$ (but possibly depending on $\tau$). By the property of multivariate Gaussian integrals,
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x$ $=\int_{\mathbb{R}^{d}}|x^{t}|^{2}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\int_{\mathbb{R}^{d}}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathrm{d}x$ (A.19)
$\leq C\int_{\mathbb{R}^{d}}|x^{t}|^{2}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})$ (A.20)
$\leq C,$ (A.21)
where the constant $C$ does not depend on $x^{1:t-1}$ due to Assumption 2-2. Therefore, we have
$I_{1}(R)\leq C\int_{|x^{1:t-1}|>R}\mathbb{P}_{1:t-1}(\mathrm{d}x^{1:t-1})=C\mathbb{P}_{1:t-1}(|X^{1:t-1}|>R)<\varepsilon_{\rm score}^{2}/4,$ (A.22)
with a sufficiently large $R$. On the other hand, for some $0<\eta<1/2$, if $|x^{t}|\leq\eta|x|$, then $|x-x^{t}h_{1}(\tau)|^{2}\geq\frac{1}{2}|x|^{2}-|h_{1}(\tau)x^{t}|^{2}\geq(\frac{1}{2}-\eta)|x|^{2}$, hence
$\int_{|x|>R}\int_{|x^{t}|\leq\eta|x|}|x|^{2}e^{-\frac{|x^{t}-xh_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x\leq$ $\int_{|x|>R/2}\int_{|x^{t}|\leq\eta|x|}|x^{t}|^{2}e^{-\frac{(\frac{1}{2}-\eta)|x|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x$ (A.23)
$\leq\bigg{(}\int_{\mathbb{R}^{d}}|x^{t}|^{2}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\bigg{)}\bigg{(}\int_{|x|>R}e^{-\frac{(\frac{1}{2}-\eta)|x|^{2}}{2h_{2}(\tau)}}\mathrm{d}x\bigg{)}$ (A.24)
$\leq CR^{d}e^{-\frac{(\frac{1}{2}-\eta)R^{2}}{2}},$ (A.25)
where we use Assumption 2-2 and the property of Gaussian integrals. Furthermore, by a similar argument as in (A.21),
$\int_{|x|>R}\int_{|x^{t}|>\eta|x|}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x\leq$ $\int_{|x|>R}\int_{|x^{t}|>\eta R}|x|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x$ (A.26)
$\leq$ $\int_{|x^{t}|>\eta R}|x^{t}|^{2}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\int_{\mathbb{R}^{d}}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathrm{d}x$ (A.27)
$\leq$ $C\mathbb{E}_{\mathbb{P}_{x^{1:t-1}}}|X^{t}|^{2}I_{\{|X^{t}|>\eta R/2\}}.$ (A.28)
Combining (A.25) and (A.28) and taking $R$ sufficiently large in the order $e^{-cR^{2}}\sim\varepsilon_{\rm score}$, we have
$I_{2}(R)=$ $\int_{\mathbb{R}^{d(t-1)}}\int_{|x|>R}\int_{\mathbb{R}^{d}}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x\mathbb{P}_{1:t-1}(\mathrm{d}x^{1:t-1})$ (A.29)
$=$ $\int_{\mathbb{R}^{d(t-1)}}\Bigg{[}\int_{|x|>R}\int_{|x^{t}|\leq\eta|x|}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}x$ (A.30)
$+\int_{|x|>R}\int_{|x^{t}|>\eta|x|}|x^{t}|^{2}e^{-\frac{|x-x^{t}h_{1}(\tau)|^{2}}{2h_{2}(\tau)}}\mathbb{P}_{x^{1:t-1}}(\mathrm{d}x^{t})\mathrm{d}\Bigg{]}\mathbb{P}_{1:t}(\mathrm{d}x^{1:t-1})$ (A.31)
$<\frac{\varepsilon_{\rm score}^{2}}{4}.$ (A.32)
From (A.18), (A.22) and (A.32) we conclude
$\mathbb{E}_{X^{1:t-1}\sim\mathbb{P}_{1:t-1}}\mathbb{E}_{X_{\tau}^{t}\sim p_{t}(\tau,\cdot|X^{1:t-1})}\Bigg{[}\bigg{|}G_{\theta}(\tau,X^{1:t-1},X^{t}_{\tau})-G(\tau,X^{1:t-1},X^{t}_{\tau})\bigg{|}^{2}I_{B_{R}^{c}}\Bigg{]}<\frac{\varepsilon_{\rm score}^{2}}{2}.$ (A.33)
From Assumption 2-1 and Lemma A.1333By an additional ReLu layer, the support set can be transferred from $0,1]^{d}$ to $B_{R}$. See Appendix B.1 of Chen et al. (2023a) for details., by choosing appropriate $\theta$, it can be achieved that
$\mathbb{E}_{X^{1:t-1}\sim\mathbb{P}_{1:t-1}}\mathbb{E}_{X_{\tau}^{t}\sim p_{t}(\tau,\cdot|X^{1:t-1})}\Bigg{[}\bigg{|}G_{\theta}(\tau,X^{1:t-1},X^{t}_{\tau})-G(\tau,X^{1:t-1},X^{t}_{\tau})\bigg{|}^{2}I_{B_{R}}\Bigg{]}<\frac{\varepsilon_{\rm score}^{2}}{2}.$ (A.34)
Based on (A.33) and (A.34) we establish (A.9). ∎
Appendix B Proof of Theorem 2.3
To obtain the Wasserstein bounds from total variation bounds, we first provide a uniform-in-time moment estimate for SDE under strong dissipative condition; see Section 1.2 of Khasminskii (2012). Suppose that for the function $f:[0,\mathcal{T}]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, there exists $M_{\rm disp},\delta>0$ such that for any $t,x$,
$x\cdot f(t,x)\leq-(1+\delta)|x|^{2}+M_{\rm disp}.$ (B.1)
We consider the following SDE:
$\mathrm{d}X_{t}=(X_{t}+f(t,X_{t}))\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$ (B.2)
$X_{0}\sim N(0,I).$ (B.3)
Lemma B.1.
There exist constants $c_{1},c_{2},c_{3}>0$, which are independent from $t$ and $M_{\rm disp}$, such that
$\sup_{t\geq 0}\mathbb{E}[e^{c_{1}|X_{t}|^{2}}]\leq c_{2}e^{c_{3}M_{\rm disp}}.$ (B.4)
Proof.
Let $V(x)=e^{\theta|x|^{2}}$, where $\theta>0$ will be determined later. Applying Itô’s formula to $V(X_{s})$, we have
$V(X_{t})-V(X_{0})=$ $\int_{0}^{t}\bigg{(}2\theta\Big{(}V(X_{s})X_{s}\cdot\big{(}X_{s}+f(s,X_{s})\big{)}\Big{)}+2\theta dV(X_{s})+4\theta^{2}|X_{s}|^{2}V(X_{s})\bigg{)}ds$ (B.5)
$+2\sqrt{2}\theta\int_{0}^{t}X_{s}V(X_{s})dB_{s}.$ (B.6)
Considering a sequence of localization stopping time $\tau_{n}\to\infty$ and taking expectation on the above identity (B.5), we obtain
$\mathbb{E}[V(X_{t\wedge\tau_{n}})]-\mathbb{E}[V(X_{0})]=$ $\int_{0}^{t\wedge\tau_{n}}\bigg{(}2\theta\Big{(}\mathbb{E}\big{[}V(X_{s})X_{s}\cdot\big{(}X_{s}+f(s,X_{s})\big{)}\big{]}\Big{)}$ (B.7)
$+2\theta d\mathbb{E}[V(X_{s})]+4\theta^{2}\mathbb{E}[|X_{s}|^{2}V(X_{s})]\bigg{)}ds$ (B.8)
$=$ $\int_{0}^{t}\bigg{(}2\theta\Big{(}\mathbb{E}\big{[}V(X_{s\wedge\tau_{n}})X_{s\wedge\tau_{n}}\cdot\big{(}X_{s\wedge\tau_{n}}+f(s\wedge\tau_{n},X_{s\wedge\tau_{n}})\big{)}\big{]}\Big{)}$ (B.9)
$+2\theta d\mathbb{E}[V(X_{s\wedge\tau_{n}})]+4\theta^{2}\mathbb{E}[|X_{s\wedge\tau_{n}}|^{2}V(X_{s\wedge\tau_{n}})]\bigg{)}ds$ (B.10)
The differential form yields
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[V(X_{t\wedge\tau_{n}})]=$ $2\theta\mathbb{E}\big{[}V(X_{s\wedge\tau_{n}})X_{s\wedge\tau_{n}}\cdot\big{(}X_{s\wedge\tau_{n}}+f(s\wedge\tau_{n},X_{s\wedge\tau_{n}})\big{)}\big{]}\Big{)}$ (B.11)
$+2\theta d\mathbb{E}[V(X_{t\wedge\tau_{n}})]+4\theta^{2}\mathbb{E}[|X_{t\wedge\tau_{n}}|^{2}V(X_{t\wedge\tau_{n}})]$ (B.12)
$\leq$ $(4\theta^{2}-2\theta\delta)\mathbb{E}[|X_{t\wedge\tau_{n}}|^{2}V(X_{t\wedge\tau_{n}})]+2\theta(M_{\rm disp}+d)\mathbb{E}[V(X_{t\wedge\tau_{n}})].$ (B.13)
Choose $\theta<\delta/2$, $R^{2}>2(M_{\rm disp}+d)/(\delta-2\theta)$ and denote $I_{t}=\big{(}(4\theta^{2}-2\theta\delta)|X_{t\wedge\tau_{n}}|^{2}+2\theta(M_{\rm disp}+d)\big{)}V(X_{t\wedge\tau_{n}})$. On the one hand, with $\gamma:=2\theta(M_{\rm disp}+d)$, it is clear that $I_{t}\leq-\gamma V(X_{t\wedge\tau_{n}})$ when $|X_{t\wedge\tau_{n}}|\geq R$. On the other hand, when $|X_{t\wedge\tau_{n}}|\leq R$,
$I_{t}\leq$ $2\theta(M_{\rm disp}+d)V(X_{t\wedge\tau_{n}})$ (B.14)
$\leq$ $2\theta(M_{\rm disp}+d)V(R)$ (B.15)
$=$ $2\theta(M_{\rm disp}+d)e^{\frac{2\theta(M_{\rm disp}+d)}{\delta-2\theta}}$ (B.16)
$\leq$ $c^{\prime}_{1}e^{c^{\prime}_{2}M_{\rm disp}},$ (B.17)
where $c^{\prime}_{1}$, $c^{\prime}_{2}$ are some constants independent of $t$ and $M_{\rm disp}$. We thus obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}\big{[}V(X_{t\wedge\tau_{n}})]\leq$ $-\gamma\mathbb{E}[V(X_{t\wedge\tau_{n}})I_{|X_{t\wedge\tau_{n}}|>R}]+c^{\prime}_{1}e^{c^{\prime}_{2}M_{\rm disp}}$ (B.18)
$=$ $-\gamma\mathbb{E}[V(X_{t\wedge\tau_{n}})]+\gamma\mathbb{E}[V(X_{t\wedge\tau_{n}})I_{|X_{t\wedge\tau_{n}}|\leq R}]+c^{\prime}_{1}e^{c^{\prime}_{2}M_{\rm disp}}$ (B.19)
$\leq$ $-\gamma\mathbb{E}[V(X_{t\wedge\tau_{n}})]+c^{\prime}_{3}e^{c^{\prime}_{4}M_{\rm disp}}$ (B.20)
By Grownwall’s inequality,
$\mathbb{E}\big{[}V(X_{t\wedge\tau_{n}})]\leq e^{-\gamma t}\mathbb{E}[V(X_{0})I_{|X_{0}|>R}]+\frac{c^{\prime}_{3}e^{c^{\prime}_{4}M_{\rm disp}}}{\gamma}\leq c^{\prime}_{5}e^{c_{6}^{\prime}M_{\rm disp}},$
Here, constants $c^{\prime}_{1}\sim c^{\prime}_{6}$ are all independent of $t$, $n$ and $M_{\rm disp}$. Thus, letting $n\to\infty$, (B.4) is established with $c_{1}=\theta$, $c_{2}=c^{\prime}_{5}$, and $c_{3}=c^{\prime}_{6}$, all independent of $t$ and $M_{\rm disp}$. ∎
Corollary B.2.
For any $t,R>0$, there exists $c_{1},c_{2}>0$, which are independent of $t$ and $R$, such that
$\mathbb{E}[|X_{t}|^{2}I_{\{|X_{t}|\geq R\}}]\leq e^{-c_{1}R^{2}+c_{2}M_{\rm disp}}.$ (B.21)
Proof of Theorem 2.3.
(1) With initial conditions $\bar{X}_{0}\sim p(\mathcal{T},\cdot|x^{1:t})$ and $\bar{Y}_{0}\sim\mathcal{N}(0,I)$, consider the following two SDEs:
$\mathrm{d}\bar{X}_{\tau}=[\bar{X}_{\tau}+2\nabla_{x}\log p_{t+1}(\mathcal{T}-\tau,\bar{X}_{\tau}|x^{1:t})]\mathrm{d}\tau+\sqrt{2}\mathrm{d}B_{\tau},$ (B.22)
$\mathrm{d}\bar{Y}_{\tau}=[\bar{Y}_{\tau}+2s^{t+1}_{\theta}(\mathcal{T}-\tau,y^{1:t},\bar{Y}_{\tau})]\mathrm{d}\tau+\sqrt{2}\mathrm{d}B_{\tau}.$ (B.23)
By the proof of total variation bounds for continuous time DDPM (see, e.g., Song et al. (2021a) and Chen et al. (2023b)), we have
$\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})^{2}\leq$ $C\bigg{(}e^{-2\mathcal{T}}+\int_{0}^{\mathcal{T}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|x^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,y^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|x^{1:t})\big{|}^{2}\mathrm{d}\tau\bigg{)}$ (B.24)
$\leq$ $C\big{(}e^{-2\mathcal{T}}+\mathcal{E}(x^{1:t})+\mathcal{T}|x^{1:t}-y^{1:t}|^{2}\big{)}$ (B.25)
Therefore, recalling that $\sqrt{a+b+c}\leq\sqrt{a}+\sqrt{b}+\sqrt{c}$ for $a,b,c\geq 0$, we have
$\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})\leq C\big{(}e^{-\mathcal{T}}+\mathcal{E}(x^{1:t})^{1/2}+\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|\big{)}.$ (B.26)
To bound $\mathcal{W}(\mathbb{P}_{x^{1:t}},\mathbb{Q}_{y^{1:t}})$ with their total variation metric, we use the following classical results regarding these two different metrics (see Villani (2008), Theorem 6.15): for any two probability measures $\mu,\nu$ and any $R>0$,
$\mathcal{W}_{2}^{2}(\mu,\nu)\leq R^{2}\mathrm{TV}(\mu,\nu)+\int_{|x|\geq R}|x|^{2}(\mu+\nu)(\mathrm{d}x).$ (B.27)
Taking $\mu=\mathbb{P}_{x^{1:t}}$, $\nu=\mathbb{Q}_{y^{1:t}}$, we have
$\mathcal{W}_{2}^{2}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})\leq C\bigg{(}R^{2}\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})+\mathbb{E}_{\mathbb{P}_{x^{1:t}}}\big{[}|X^{t+1}|^{2}I_{\{|X^{t+1}|\geq R\}}\big{]}+\mathbb{E}_{\mathbb{Q}^{\mathcal{T}}_{y^{1:t}}}\big{[}|Y^{t+1}|^{2}I_{\{|Y^{t+1}|\geq R\}}\big{]}\bigg{)}.$ (B.28)
We now estimate three terms on the right-hand side of (B.28). By Assumption 2,
$\mathbb{E}_{\mathbb{P}_{x^{1:t}}}\big{[}|X^{t+1}|^{2}I_{\{|X^{t+1}|\geq R\}}\big{]}\leq$ $C\mathbb{E}_{\mathbb{P}_{x^{1:t}}}\big{[}e^{\frac{c}{2}|X^{t+1}|}I_{\{|X^{t+1}|\geq R\}}\big{]}$ (B.29)
$\leq$ $C\Big{\{}\mathbb{E}_{\mathbb{P}_{x^{1:t}}}\big{[}e^{c|X^{t+1}|}\big{]}\Big{\}}^{1/2}\Big{\{}\mathbb{P}_{x^{1:t}}\big{(}|X^{t+1}|\geq R\big{)}\Big{\}}^{1/2}$ (B.30)
$\leq$ $Ce^{-\frac{c}{2}R}\mathbb{E}_{\mathbb{P}_{x^{1:t}}}\big{[}e^{c|X^{t+1}|}\big{]}.$ (B.31)
$\leq$ $Ce^{-\frac{c}{2}R}$ (B.32)
By Corollary B.2,
$\mathbb{E}_{\mathbb{Q}_{y^{1:t}}}\big{[}|Y^{t+1}|^{2}I_{\{|Y^{t+1}|\geq R\}}\big{]}$ $=\mathbb{E}_{\mathbb{Q}^{\mathcal{T}}}[|\bar{Y}_{\mathcal{T}}|^{2}I_{\{|\bar{Y}_{\mathcal{T}}|\geq R\}}\big{]}$ (B.33)
$\leq e^{-c_{1}R^{2}+c_{2}M_{\rm disp}}.$ (B.34)
Plugging (B.26), (B.32) and (B.33) into (B.28), we have
$\mathcal{W}_{2}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})^{2}\leq$ $C(R^{2}e^{-\mathcal{T}}+R^{2}\mathcal{E}(x^{1:t})^{1/2}+R^{2}\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|+Ce^{-\frac{c}{2}R}+e^{-c_{1}R^{2}+c_{2}M_{\rm disp}})$ (B.35)
$\leq$ $C(R^{2}e^{-\mathcal{T}}+R^{2}\mathcal{E}(x^{1:t})^{1/2}+R^{2}\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|+e^{-cR+cM_{\rm disp}}).$ (B.36)
Taking $R=\mathcal{T}$, we conclude (2.31).
(2). Because $\pi\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})$ if and only if $\pi_{x^{1:t},y^{1:t}}\in\Pi(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})$ (Backhoff et al. (2017), Proposition 5.1), we construct a specific coupling $\pi^{\varepsilon}\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ as follows:
  • (a)
    $\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x^{1}-y^{1}|^{2}\pi^{\varepsilon}_{1}(\mathrm{d}x^{1},\mathrm{d}y^{1})\leq\mathcal{W}_{2}^{2}(\mathbb{P}_{1},\mathbb{Q}^{\mathcal{T}}_{1})+\varepsilon$;
  • (b)
    For $t=1,2,\cdots,T-1$, $\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x^{t+1}-y^{t+1}|^{2}\pi^{\varepsilon}_{x^{1:t},y^{1:t}}(\mathrm{d}x^{t+1},\mathrm{d}y^{t+1})\leq\mathcal{W}_{2}^{2}(\mathbb{P}_{x^{1:t}},\mathbb{Q}_{y^{1:t}})+\varepsilon$;
  • (c)
    $\pi^{\varepsilon}(\mathrm{d}x^{1:T},\mathrm{d}y^{1:T}):=\pi^{\varepsilon}_{1}(\mathrm{d}x^{1},\mathrm{d}y^{1})\pi^{\varepsilon}_{x^{1},y^{1}}(\mathrm{d}x^{2},\mathrm{d}y^{2})\cdots\pi^{\varepsilon}_{x^{1:T-1},y^{1:T-1}}(\mathrm{d}x^{T},\mathrm{d}y^{T})$.
To prove (2.35), we first use induction to show that, for $t=1,2,\cdots,T$,
$\mathbb{E}_{\pi^{\varepsilon}_{1:t}}|X^{1:t}-Y^{1:t}|^{2}\leq C\Big{(}\mathcal{T}^{\frac{5t}{2}}\varepsilon_{\rm score}^{1/2^{t-1}}+\mathcal{T}^{\frac{5(t-1)}{2}}(\alpha(\mathcal{T})+\varepsilon)^{1/2^{t-1}}\Big{)}.$ (B.37)
By repeating the proof of (1), with $\mathbb{P}_{x^{1:t}}$, $\mathbb{Q}_{y^{1:t}}$ replaced by $\mathbb{P}_{1}$, $\mathbb{Q}_{1}$, we can prove that (B.37) holds for $t=1$. This is because when considering non-conditional probability $\mathbb{P}_{1}$ and $\mathbb{Q}_{1}$, the term $|x^{1:t}-y^{1:t}|$ disappears and the term $\mathcal{E}(x^{1:t})$ is upper bounded by $\mathcal{T}\varepsilon_{\rm score}^{2}$.
Suppose now that (B.37) is true for $t$, we aim to prove it for $t+1$. Indeed, by using (2.31), we have
$\mathbb{E}_{\pi_{1:t+1}^{\varepsilon}}[|X^{1:t+1}-Y^{1:t+1}|^{2}]\leq$ $C\bigg{(}\mathbb{E}_{\pi^{\varepsilon}_{1:t}}[|X^{1:t}-Y^{1:t}|^{2}]+\mathbb{E}_{\pi^{\varepsilon}_{1:t}}\mathbb{E}_{\pi_{X^{1:t},Y^{1:t}}^{\varepsilon}}[|X^{t+1}-Y^{t+1}|^{2}]\bigg{)}$ (B.38)
$\leq$ $C\bigg{(}\mathbb{E}_{\pi^{\varepsilon}_{1:t}}[|X^{1:t}-Y^{1:t}|^{2}]+\mathbb{E}_{\pi^{\varepsilon}_{1:t}}\big{[}\mathcal{W}^{2}_{2}\big{(}\mathbb{P}_{X^{1:t}},\mathbb{Q}^{\mathcal{T}}_{Y^{1:t}}\big{)}\big{]}+\varepsilon\bigg{)}$ (B.39)
$\leq$ $C\bigg{(}\mathbb{E}_{\pi^{\varepsilon}_{1:t}}|X^{1:t}-Y^{1:t}|^{2}+\mathcal{T}^{2}\mathbb{E}_{\pi^{\varepsilon}_{1:t}}\mathcal{E}(X^{1:t})^{1/2}+\mathcal{T}^{\frac{5}{2}}\big{\{}\mathbb{E}_{\pi^{\varepsilon}_{1:t}}|X^{1:t}-Y^{1:t}|^{2}\big{\}}^{1/2}$ (B.40)
$+\alpha(\mathcal{T})+\varepsilon\bigg{)}$ (B.41)
With (2.4), we have
$\mathbb{E}_{\pi^{\varepsilon}_{1:t}}[\mathcal{E}(X^{1:t})^{1/2}]=\mathbb{E}_{\mathbb{P}_{1:t}}[\mathcal{E}(X^{1:t})^{1/2}]\leq\big{\{}\mathbb{E}_{\mathbb{P}_{1:t}}[\mathcal{E}(X^{1:t})]\big{\}}^{1/2}\leq\mathcal{T}^{1/2}\varepsilon_{\rm score}.$ (B.42)
If $\mathbb{E}_{\pi^{\varepsilon}_{1:t}}|X^{1:t}-Y^{1:t}|^{2}>1$, then from (B.41) we know
$\mathbb{E}_{\pi_{1:t+1}^{\varepsilon}}|X^{1:t+1}-Y^{1:t+1}|^{2}\leq$ $C\Big{(}\mathcal{T}^{\frac{5}{2}}\mathbb{E}_{\pi_{1:t}^{\varepsilon}}|X^{1:t}-Y^{1:t}|^{2}+\mathcal{T}^{\frac{5}{2}}\varepsilon_{\rm score}+\alpha(\mathcal{T})+\varepsilon\Big{)}$ (B.43)
$\leq$ $C\Big{(}\mathcal{T}^{\frac{5(t+1)}{2}}\varepsilon_{\rm score}^{1/2^{t-1}}+\mathcal{T}^{\frac{5t}{2}}(\alpha(\mathcal{T})+\varepsilon)^{1/2^{t-1}}$ (B.44)
$+\mathcal{T}^{\frac{5}{2}}\varepsilon_{\rm score}+\alpha(\mathcal{T})+\varepsilon\Big{)}$ (B.45)
$\leq$ $C\Big{(}\mathcal{T}^{\frac{5(t+1)}{2}}\varepsilon_{\rm score}^{1/2^{t}}+\mathcal{T}^{\frac{5t}{2}}(\alpha(\mathcal{T})+\varepsilon)^{1/2^{t}}\Big{)}.$ (B.46)
Here, we use $\mathcal{T}>1$, $\varepsilon_{\rm score}<1$ and choose a small $\varepsilon$ such that $\alpha(\mathcal{T})+\varepsilon<1$. Similarly, if $\mathbb{E}_{\pi^{\varepsilon}_{1:t}}|X^{1:t}-Y^{1:t}|^{2}\leq 1$,
$\mathbb{E}_{\pi_{1:t+1}^{\varepsilon}}|X^{1:t+1}-Y^{1:t+1}|^{2}\leq$ $C\Big{(}\mathcal{T}^{\frac{5}{2}}\big{\{}\mathbb{E}_{\pi_{1:t}^{\varepsilon}}|X^{1:t}-Y^{1:t}|^{2}\big{\}}^{1/2}+\mathcal{T}^{\frac{5}{2}}\varepsilon_{\rm score}+\alpha(\mathcal{T})+\varepsilon\Big{)}$ (B.47)
$\leq$ $C\Big{(}\mathcal{T}^{\frac{5}{2}\big{(}\frac{t}{2}+1\big{)}}\varepsilon_{\rm score}^{1/2^{t}}+\mathcal{T}^{\frac{5}{2}\big{(}\frac{t-1}{2}+1\big{)}}(\alpha(\mathcal{T})+\varepsilon)^{1/2^{t}}$ (B.48)
$+\mathcal{T}^{\frac{5}{2}}\varepsilon_{\rm score}+\alpha(\mathcal{T})+\varepsilon\Big{)}$ (B.49)
$\leq$ $C\Big{(}\mathcal{T}^{\frac{5(t+1)}{2}}\varepsilon_{\rm score}^{1/2^{t}}+\mathcal{T}^{\frac{5t}{2}}(\alpha(\mathcal{T})+\varepsilon)^{1/2^{t}}\Big{)}.$ (B.50)
In either case, we establish that (B.37) holds for $t+1$, hence for all $t\in\{1,2,\cdots,T\}$. Taking $t=T$ in (B.37) and using the definition of adapted Wasserstein metric, we obtain
$\mathcal{A}\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq$ $\mathbb{E}_{\pi^{\varepsilon}}|X^{1:T}-Y^{1:T}|^{2}$ (B.51)
$\leq$ $C\Big{(}\mathcal{T}^{\frac{5T}{2}}\varepsilon_{\rm score}^{1/2^{T-1}}+\mathcal{T}^{\frac{5(T-1)}{2}}(\alpha(\mathcal{T})+\varepsilon)^{1/2^{T-1}}\Big{)}.$ (B.52)
Letting $\varepsilon\to 0$ in (B.52) gives (2.35). ∎